### v1

Abstract
[…]
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
*Equal contribution
1 Introduction
Today, large-language model (LLM) inference has outpaced training in growth. LLM inference powers a wide range of applications, from interactive customer support and code generation to retrieval-based document analysis and agentic workflows. However, the LLM inference systems have not caught up—latency (both response latency and generation speed) and cost are now the bottleneck.
A key limitation of today’s LLM inference systems is that each user query is independently processed by one instance of the inference engine, without any data being reused across queries or inference engines. This means that a LLM query’s lifecycle, including computation and I/O operations, happens in the GPUs and the GPU memory of one inference engine. Each query takes a token sequence as input and another token sequence as output, and once a query completes, all outputs or intermediate states generated during inference are discarded.
Figure 1: LMCache supports both context caching (KV cache offloading and sharing across queries) and PD disaggregation (cross-engine transfer of KV caches.
This design, albeit simple, leads to redundant computation and resource underutilization, and two notable industry proposals attempted to address them (Figure 1).
Cross-query caching to avoid redundant compute:  As each query is treated in isolation, the LLM must prefill the same context (prefix) each time the context is used, and the repeated prefill compute is expensive and causes slow response, as users cannot see the first token before the prefill completes. Yet, such redundant computation could be avoided if we persist KV cache of the prefix—which LLM created internally after it prefills the prefix—beyond the lifecycle of a query and reuse the KV cache when subsequent queries reuse the prefix (e.g., the same prompt template or retrieved document chunk). By avoiding redundant computation, cross-query context (prefix) caching significantly lowers both time-to-first-token (TTFT) and overall GPU resource consumption during the prefill phase.
Prefill–decode disaggregation for higher utilization:  Co-locating the prefill phase (compute-bound, throughput-oriented) and the decode phase (memory-bound, latency-sensitive) in the same inference engine causes resource underutilization as the GPUs must be overprovisioned to ensure a consistently low decoding latency without being interrupted by prefill. In response, the industry attempted to decouple the throughput-oriented prefill processing from latency-sensitive decoding Such prefill-decode (PD) diaggregation architecture batches all queries’ prefill phase on one set of inference nodes and transfers their KV caches to another set of inference nodes that run the decoding phase only, thus mitigating tail decoding latency under high concurrency.
To make these ideas practical, the LLM inference systems must be augmented with new KV cache semantics. In particular, inference engines should support the new interface that extracts KV caches from a normal inference call and re-loads KV caches into subsequent queries on demand. The system also must allow the extracted KV caches to be stored persistently and transferred across distributed inference engines. Most importantly, for such interface extensions to be practical, KV cache extraction, re-loading, storage, and transfer must be efficient, and the new interfaces must remain compatible with rapidly evolving inference engines such as vLLM Kwon et al. (2023b) and SGLang Zheng et al. (2024).
We introduce LMCache, the first open-source library that provides a high-performance implementation of these new KV cache semantics. With LMCache, KV cache can be extracted from and loaded back to inference engines efficiently, stored in a hierarchy of storage devices (CPU memory, local disk, remote disk, and Redis), and transferred over different networks (Ethernet, RDMA, NVLink).
LMCache makes three distinct contributions.
#1. Highly optimized performance:  LMCache incorporates a series of performance optimizations that make storing and loading KV cache efficient and practical in real deployments. For instance, LMCache batches operations to pipeline the storing and loading of KV cache, as well as to pipeline GPU compute and data loading/storing (e.g., loading next layer’s KV cache while performing computations for the current layer). Moreover, rather than storing/loading KV cache at the granularity of the inference engine’s native small page size, LMCache stores/loads KV cache at a configurable chunk size, often much larger than page size, to fully utilize the bandwidth between storage devices and GPU memory. LMCache also minimizes the copies of KV cache data when moving them among different storage tiers, by implementing zero-copy operations.
#2. Standardized interface with inference engines:  LMCache defines standardized connector interfaces that remain compatible with fast-changing inference engine backends. On average, 15–20 new open-weight models are released every week in 2025, so to best utilize new hardware for the new models, modern LLM inference engines must evolve rapidly, potentially changing the KV cache layout in GPU memory and thus affecting the LMCache interface. To address this, LMCache designs and implements a modular KV cache connector interface that decouples LMCache with the inference engine backend, so LMCache can easily adapt to the evolving APIs in the inference engines.
#3. Flexible KV cache management interface:  The interface augmentation introduced by LMCache exposes KV cache, a new data structure in LLM inference. LMCache exposes APIs that allow developers and operators to locate, move, pin, and even compress KV cache extracted from inference engines. These first-class APIs allow higher-level applications, such as query schedulers or routers, to make better decisions, such as KV cache-aware query routing.
[…]
The remainder of this paper details the motivation (§2), LMCache architecture and key design choices (§3–§6), experimental evaluation (§8), and deployment experiences (§9).
2 Motivation
2.1 Inference Cost and Delay as the Bottleneck
The past two years have witnessed explosive growth in the deployment of LLMs across diverse enterprise applications. As the user base and query volume of LLM-based applications increase, the cost and delay of inference, rather than training, emerges as the primary bottleneck to scaling LLM serving to production scale.
Inference cost:  Besides the simple fact that LLMs are queried by more users, the increasing input length and output length also contribute to raising inference cost, as each input token and each output token incur an additional cost to process. Four trends are behind the increase in input and output lengths. (i) More interactions by each user with LLMs build a longer user history, which will be prepended to future LLM inputs as the context. (ii) LLMs increasingly take multimodal inputs, such as images and videos, which typically will be transformed to a long sequence of tokens before being fed to LLMs. (iii) Newer LLMs often have longer context windows, allowing context/prompt engineers to simply stuff more tokens into LLM inputs. (iv) Finally, reasoning models produce unusually long output, which might be used as input to subsequent queries in one workflow.
Even with steady improvement on GPU price and performance, serving a trillion-token-scale workload can incur millions of dollars in annual compute expenditure Nie et al. (2024); Databricks Research (2024). Redundant computation across queries (e.g., recomputing the same prompt prefix hundreds of times) exacerbates these costs Zhang et al. (2024); Caylent (2024).
Inference delay:  In interactive applications, tail latency (e.g., 95th/99th percentile) is as critical as median latency. Slow responses in the worst case, often due to resource contention or large context processing, directly degrade user experience and can lead to user abandonment, in much the same way video streaming stalls affect viewer experience Liu et al. (2024a).
In particular, empirical observations from industry reports and academic studies indicate that:
  • $\bullet$
    Time-to-first-token (TTFT) can be dominated by the prefill phase for long-context inputs, especially when context lengths exceed 8–16k tokens.
  • $\bullet$
    Inter-token latency (ITL) remains low for single-stream decoders, but rises sharply when GPU compute is shared across many concurrent sessions.
  • $\bullet$
    Tail delays are disproportionately impacted by scenarios requiring large prompt recomputation or reallocation of memory-bound resources.
2.2 KV Caching as a Cross-Cutting Optimization
KV cache was originally introduced to accelerate a single inference query by storing the attention states, in the form of $K$ and $V$ tensors, for input tokens and previously generated tokens directly in GPU memory. KV cache effectively stores the attention information between each pair of tokens that have been seen so far in this query. In short, it is a LLM-native representation of knowledge.
In all popular transformer architectures, these caches are stored and used only within the lifecycle of an individual query and discarded when the query ends. Yet, KV caching can be extended to support two critical performance optimizations in production-scale serving:
  • 1.
    Context caching (i.e.,cross-query KV cache reuse) persists KV cache segments from one query and reusing them for subsequent queries that share a common prefix. Examples include document analysis where the same document (chunk) remains constant across multiple queries, and multi-turn dialogues with a fixed system prompt or long preamble. Prefix caching reduces redundant computation during the prefill phase, directly lowering TTFT and GPU-hours per query Liu et al. (2024b); Gao et al. (2024); Jin et al. (2025b); Ren et al. (2025); Qin et al. (2025a); Jin et al. (2024); Chen et al. (2024; 2025).
  • 2.
    Prefill–decode (PD) disaggregation (i.e.,cross-engine KV cache transfer) splits inference into a prefill stage (processing the entire input prompt) and a decode stage (autoregrssive token generation) across different GPUs or nodes. This approach reduces tail latency by maximizing the decoding speed without being interrupted by the prefill phase Zhong et al. (2024); Patel et al. (2024); Shi et al. (2025).
Roles of KV caching:  Both optimizations critically depend on the ability to move KV cache segments efficiently to and from GPU memory, CPU DRAM, NVMe storage, or over the network.
  • $\bullet$
    Context reusing stores the KV cache of a context for as long as the context will be reused, so it requires a hierarchical storage of KV caches and efficient sharing of KV caches across multiple serving engine instances.
  • $\bullet$
    PD disaggregation requires fast KV cache transfer between GPUs, often across PCIe, NVLink, or RDMA, in order to minimize the end-to-end delay of each query, which runs across the prefiller nodes and the decoder nodes.
2.3 Challenges of Efficient KV Caching
Despite their potential, the practical adoption of prefix caching and PD disaggregation is limited by three interrelated systems challenges:
Message Size Transfer Throughput
64KB 4GBps
256KB 13GBps
1MB 30GBps
10MB 46GBps
16MB 49GBps
100MB 49GBps
Table 1: Transfer message size vs achieved transfer throughput using RCCL transfer library UCCL Team (2025).
2.3.1 Challenge #1: I/O inefficiency under paged memory
KV cache storage and transfer used to rely on PyTorch serialization (torch.save / torch.load) or primitive tensor copying, with a typical transfer speed of only sub-1GB/s. These methods introduce non-trivial delay overhead, especially when handling large data structures like KV caches, and lack zero-copy support with various storage devices (local or remote), causing extra CPU-GPU data copies.
Recent high-throughput inference engines, such as vLLM Kwon et al. (2023b) and SGLang Zheng et al. (2024), make KV cache storage and transfer even more challenging. They employ paged attention memory, dividing the attention buffer into small, fixed-size pages (typically 16–64 KB). For instance, vLLM uses 62.5-KB page in Llama-3.1-8B-Instruct model. The paged memory architecture is widely used because it improves batching and memory utilization.
However, because the pages of a KV cache are not always contiguous, the paged memory architecture dramatically increases the number of small-sized I/O operations required to persist or transfer a KV cache. Transferring such small chunks of data is known to suffer from network bandwidth underutilization and reduce throughput Kwon et al. (2025); NVIDIA Developer Forums (2020); Meta Engineering (2024). Prior work (Table 1) has shown that, on a setup with two AMD GPU nodes connected by eight Broadcom Thor-2 400Gbps NICs, the transfer size must reach at least 16 MB to saturate the available network bandwidth Zhou et al. (2025). Furthermore, prior work has shown that only transferring a data size of megabyte range (e.g., 1–2MB) can achieve 75–80% of the theoretical PCIe 5.0 bandwidth Xie et al. (2025).
2.3.2 Challenge #2: Compatible with fast-evolving inference engines
With the widespread use of AI, new LLMs and hardware accelerators are introduced at a rapid pace. In 2025, one prominent LLM was released on average every 4 days bes (2025). In response, inference engines must evolve just as quickly.
Each update to accommodate new models or hardware often changes GPU memory allocation, which in turn changes the KV cache interface. For example, when vLLM adopts a new attention kernel that produces KV caches with different dimensions, the KV caching library must be updated to translate the new kernel’s output KV cache format into one compatible with the KV cache library. Keeping up with these frequent changes requires tremendous effort, given the fast-moving inference engines.
2.3.3 Challenge #3: Lack of management APIs
As KV caching becomes a first-class citizen in the LLM inference backend, various components (in addition to the LLM inference engines), as well as ML ops teams, will need to make decisions in a KV-cache-aware manner. Yet, without a unified management interface to locate, evict, pin, or compress caches, these upper-layer modules cannot make informed placement or eviction decisions. This leads to inefficient cache utilization, duplicated storage, and unpredictable eviction policies. For instance, inference query routers, which assign each query to one of the inference engine instances, need to know the locations of KV caches, in order to route queries to instances that already hold the KV cache for matched prefix tokens locally (e.g., in CPU memory).
Moreover, applications now also demand such KV-cache management interfaces. In early 2025, for instance, a financial company [^1] For confidentiality, we do not disclose names of enterprise users in this report. that has worked closely with LMCache in the production setting asked for an interface that allows users to explicitly pin frequently accessed financial documents in the KV caching system, for more efficient access to popular contexts. As another example, an agent company requested a series of APIs that allow them to identify the KV cache of a given content, compress the KV cache, and transfer the compressed KV cache across nodes.
2.4 Related Work and Existing Solutions
Several KV cache handling mechanisms exist, but none of them fully address the above challenges:
Inference frameworks:  Since the release of vLLM Production Stack vLLM project (2025) in January 2025, there have been several open-source distributed inference stacks, including Nvidia’s Dynamo NVIDIA Corporation (2025), AIBrix Team et al. (2025), llm-d llm-d Project (2025), SGLang OME Team (2025), and KServe Contributors (2025b). They focus on easy deployment of inference engine solutions over Kubernetes, and technically, they all support various query routers based on load or prefix cache awareness and support KV caching, where LMCache is used in vLLM production stack, Dynamo, llm-d, and KServe.
Inference engine-native KV caching:  Open-source inference engines, like vLLM and SGLang, also offer native GPU-to-CPU KV cache transfers, but it is designed for single-node inference, so they lack cross-node transfer optimization or hierarchical storage support for KV cache. We will evaluate their performance and compare it against LMCache in §7.
KV cache storage layers:  Mooncake Qin et al. (2025b), Redis Redis (2025), InfiniStore ByteDance (2025), and 3FS Contributors (2025a) provide distributed object storage or caching, but they lack an efficient “glue” layer between the inference engines to efficiently move small tensors frequently across different storage tiers, or are tightly coupled with a specific inference framework.
Proprietary implementations:  Proprietary inference APIs (e.g., Fireworks AI, Together AI) implement their own prefix caching internally, but these are tied to their closed-source serving stacks and are not accessible to operators deploying their own infrastructure.
Source code for research:  Several research proposals have open-sourced prototypes for their KV cache optimizations, including prefix caching Kwon et al. (2023a); Zheng et al. (2024); Yu et al. (2025); Gim et al. (2024); Ye et al. (2024); Lee et al. (2024); Zhao et al. (2024); Jin et al. (2024); Gao et al. (2024); Chen et al. (2025); Jin et al. (2025a); Yang et al. (2025), PD disaggregation Zhong et al. (2024); Patel et al. (2024); Shi et al. (2025), and KV cache compression Liu et al. (2024c); Jegou et al. (2024); Xiao et al. (2024a; b); Li et al. (2024); Qin et al. (2025c); Tang et al. (2024); Ge et al. (2024); Li et al. (2025); Du et al. (2025); Zhang et al. (2025). However, these prototypes are typically built on research-oriented inference framworks, such as HuggingFace Transformers, not fully enterprise-ready, or are not designed to evolve alongside the rapidly changing inference engine ecosystem, such as SGLang and vLLM.
3 Overview of LMCache
LMCache addresses these challenges by a unified, high-performance KV caching layer capable of efficient storage, movement, and explicit management of KV caches for paged-memory inference engines, making prefix caching and PD disaggregation practical at enterprise scale.
As a KV caching layer, LMCache sits between LLM inference engines and heterogeneous storage/network devices (Figure 2). Its goal is to provide a standardized, high-performance substrate for KV cache movement and management, while remaining compatible with rapidly evolving inference frameworks such as vLLM and SGLang.
Figure 2: LMCache sits between LLM inference engines and heterogeneous storage/network devices.
3.1 Architecture
At a high level, LMCache connects to inference engines to move KV cache between GPU memory and the storage backend. Figure 3 shows the end-to-end system. Below, we walk through two example workflows: storing and retrieving KV cache.
Store:  When a new query arrives, it first passes through the KV connector, which prepares metadata such as the tokenized input prompt and GPU memory addresses of the relevant pages. The query then goes to the token processor, which determines how many new tokens are not yet in the backend and need to be stored. Finally, the storage manager saves the KV cache for these new tokens to the backend via the transfer channel, which handles the data transfer logic.
Retrieve:  When a query requires loading KV cache from the backend, it also starts with the KV connector to prepare metadata. The token processor identifies the number of prefix-matched tokens already in the backend. Next, the event manager checks if the same query ID has been seen before. If so, the cached memory addresses are already tracked and can be returned directly to the GPU connector, which loads the KV cache back into GPU memory. The event manager also launches asynchronous, layer-wise loading events as described in §4.2. If the query ID is new, it is forwarded to the storage manager to look up the CPU memory addresses of the stored KV cache.
Lookup:  When a query needs to check whether the KV cache for specific tokens exists in the backend, higher-level components such as routers query the cache controller. The cache controller maintains a token pool that records all tokens currently stored in the KV cache backend. Whenever a LMCache instance stores or evicts a KV cache, the LMCache worker inside the instance updates the token pool with the new status. This ensures the token pool always has the up-to-date information of tokens in the backend.
Figure 3: End-to-end system workflow for LMCache.
3.2 Key Components
LMCache is composed of a high-performance data plane and a flexible control plane.
LMCache Worker (Data Plane).:  Each inference engine is paired with a LMCache worker. The LMCache worker module is responsible for moving KV caches between GPU memory and other tiers or other workers. It implements both KV offloading (to CPU or disk) and PD disaggregation (GPU–GPU transfer). To maximize throughput, workers employ kernel-optimized GPU buffers, asynchronous chunked I/O, and layer-wise pipelining, as elaborated in the next section. These techniques allow LMCache to sustain near-GPU-resident bandwidth even when handling small paged memory objects (16–64KB), a common configuration in vLLM and SGLang.
LMCache Controller (Control Plane).:  The controller exposes a programmable API to system operators and higher-level schedulers. It provides cache lifecycle management primitives, such as pinning KV segments, compressing and decompressing cache tensors, migrating caches between devices, and evicting low-priority entries. The controller also maintains a virtualized namespace, enabling uniform addressing of caches across heterogeneous devices.
Together, the LMCache worker and controller form a modular system that integrates seamlessly into enterprise-scale serving pipelines. The worker ensures fast data movement, while the controller provides higher-level control semantics necessary for efficient resource utilization.
4 Performance Optimizations
An important aspect of LMCache is improving the efficiency of KV cache movement across devices. In enterprise-scale LLM inference, LMCache addresses three key challenges:
  • $\bullet$
    modern LLM inference engines manage KV cache at the granularity of pages [^2] Each page is 16 tokens for a single layer in vLLM. , which are typically 20 KB–63 KB for popular models including Llama, Qwen, GPT-OSS etc. Such small units are inefficient for transferring, as they cannot saturate bandwidth Xie et al. (2025); Zhou et al. (2025);
  • $\bullet$
    KV cache transfers often need to run concurrently with LLM inference. This introduces overhead from two sources. First, data movement can stall inference if transfers are executed in the same CUDA stream as computation. Second, launching memory-copy CUDA functions incurs CPU overhead, as each call consumes CPU cycles and the consumption can be substantial when there are many layers and pages.
  • $\bullet$
    During LLM inference, large volumes of queries generate significant amount of KV caches. Duplicating them on any storage device wastes space and introduces copy overhead, which slows down inference.
Each of these challenges arose from hard lessons in both open-source and enterprise deployments. This section describes these challenges in detail and motivates LMCache’s design decisions.
4.1 Batched Operations
To address the I/O inefficiency caused by small KV cache units, LMCache introduces a set of optimizations.
Configurable Chunk Size:  Rather than transferring KV cache at the page level, LMCache groups multiple pages from multiple layers into larger chunks, with a default size of 256 tokens per chunk [^3] The chunk size is configurable to different I/O speeds. . This is achieved using an intermediate GPU buffer. For storing, multiple pages (16 by default) are first copied from GPU memory into the buffer, then offloaded collectively to lower-tier storage (e.g., CPU memory) and saved at the granularity of chunks rather than individual pages. For loading, chunks are first retrieved from the storage layer into the GPU buffer and subsequently split into pages before being placed in the serving engine’s internal GPU memory. LMCache implements a customized CUDA kernel to accelerate the memory copy.
Batched Store/load Operations:  LMCache supports storing and loading KV cache across multiple storage devices, including local CPU memory, local disks, and remote disks. In practice, KV caches often need to be transferred to different devices in parallel – for example, storing the KV cache of a hot context in CPU memory while simultaneously offloading a cold context to local disk. A naive approach that stores the two KV caches sequentially underutilizes available bandwidth, since transferring to CPU memory leaves the disk bandwidth idle. To address this, LMCache batches storage and loading queries across different tiers, thereby aggregating operations and fully utilizing the bandwidth between the GPU and multiple storage tiers.
Delayed Decode KV Cache Storing:  LMCache also supports saving newly generated KV cache during decoding. In a naive design, each newly generated page is immediately offloaded to lower-tier storage, incurring frequent small writes. Instead, LMCache aggregates multiple pages and stores them together as a chunk, improving I/O efficiency.
4.2 Compute-I/O Overlapping
LMCache’s data movement often happens in parallel to LLM inference computation. For example, during decoding, a KV cache is computed while being stored to storage devices at the same time. Naive data movement operations can block the LLM inference computations, or vice versa. LMCache employs multiple optimizations aiming for overlapping LLM inference computations with I/O.
Layer-wise pipelining:  LMCache overlaps KV cache transfers with inference computation through layer-wise pipelining. Specifically, it assigns separate CUDA streams for inference computation and data movement within each layer. For example, before performing inference on the first layer, its KV cache is loaded into the GPU buffer and transformed into pages. While the first layer is running inference, the KV cache for the second layer is asynchronously fetched into the buffer and similarly transformed. Note that the second layer’s KV cache loading is happened after the first layer’s KV cache is put into the right paged memory. This design ensures that only a fixed-size GPU buffer—whose size is a single layer’s KV cache—is required, while enabling overlapping between data transfer and computation.
Asynchronous compute & prefetch:  In many scenarios, there is a time gap between when the inference scheduler admits a query and when the query’s KV cache is actually needed for inference. For example, if 100 queries arrive simultaneously but the inference engine can only process 50 in parallel, the remaining 50 must wait in the queue before their inference begins. LMCache exploits this idle interval to prefetch the queued queries’ KV cache from slower storage tiers into faster ones (e.g., from remote disk to local CPU memory). As a result, when the actual inference computation starts, the required KV cache can be loaded directly from faster-tier storage, significantly reducing loading delay. LMCache allows users to configure the target tier for prefetching based on their own needs in latency SLO and resource constraints.
Process separation:  Co-locating LMCache ’s data movement with the inference engine in a single process introduces 5%–10% latency overhead due to CPU resource contention. To eliminate this overhead, LMCache decouples data movement into a separate process, isolating it from the inference computation process. Additionally, process separation also enables memory sharing across multiple inference engine instances. When data movement is co-located with computation, each inference instance maintains its own CPU memory pool, so instances cannot access one another’s KV cache without inter-process communications. By contrast, a standalone data movement process manages a unified CPU memory pool, allowing multiple inference instances to store and load KV cache from the same pool, thereby improving efficiency and reducing redundancy.
4.3 Minimum Data Copy
A naive implementation of KV cache movement would create additional copies of data at each transfer step, especially when dealing with heterogeneous storage types, leading to redundant memory usage and unnecessary overhead. LMCache avoids this by maintaining only the minimum required copies.
Zero-Copy Operations:  When transferring KV cache to remote storage or another GPU instance (e.g., in disaggregated prefill), LMCache minimizes data duplication through a reference counter. Specifically, when KV cache is written to multiple destinations—such as local CPU memory, local disk, or remote disk—LMCache increments a reference counter for the shared data instead of creating new copies. Each completed write decrements the counter, and once the count reaches zero, the data is released. This design ensures that data is shared across concurrent write operations without unnecessary replication, thus reducing memory pressure and improving efficiency. This technique is simlar to PCB counter in operating systems Strecker (1978).
Figure 4: Illustration of dynamic offloading in LMCache.
Dynamic Offloading:  Modern inference engines such as vLLM maintain a pool of free pages in GPU memory, i.e.,, pages whose KV cache is not currently used by active queries. Instead of duplicating all free pages to CPU memory, LMCache duplicates only a subset. This mechanism is implemented using three pointers:
  • $\bullet$
    Start pointer: the start address of the free-page region in GPU memory.
  • $\bullet$
    Current pointer: the index of the free pages that have already been offloaded to CPU memory.
  • $\bullet$
    End pointer: the end address of the free pages that are scheduled to be offloaded.
Function name Description
get_num_new_matched_tokens(query) $\rightarrow$ matched_tokens Returns the number of cache-hit tokens found in LMCache ’s backend.
update_state_after_alloc(query, blocks, num_external_blocks) Updates whether a query needs to transfer KV cache from LMCache’s backend.
build_connector_meta(scheduler_output) $\rightarrow$ kv_connector_metadata Builds metadata for KV cache transfers between LMCache’s backend and GPU memory, including GPU memory addresses for KV cache pages.
start_load_kv(kv_pointers) Starts loading KV cache from lower-tier storage into GPU memory before LLM inference begins.
wait_load_kv(kv_pointers, layer_id) Synchronizes on KV cache loading to ensure data is available when computation requires it.
start_store_kv(kv_pointer) Starts offloading KV cache to lower-tier storage after computation.
wait_store_kv(kv_pointer, layer_id) Synchronizes on KV cache storing to ensure the KV cache for the current layer is offloaded.
Table 2: Functions in LMCache’s connector.
As illustrated in Figure 4, dynamic offloading has four possible states:
  • $\bullet$
    State #1 (Initialization): the start and current pointers overlap. The region between the start/current pointers and the end pointer marks the pages pending duplication.
  • $\bullet$
    State #2 (In-progress): the current pointer moves toward the end pointer. Pages between the start and current pointers have already been offloaded to CPU memory.
  • $\bullet$
    State #3 (Query Arrival): when new queries acquire some of the free pages, the end pointer is moved forward by the number of allocated pages. This ensures sufficient GPU memory is available for future active queries that need to acquire free pages.
  • $\bullet$
    State #4 (Steady state): the current pointer overlaps with the end pointer, indicating that all scheduled pages have been duplicated.
Note that if a query attempts to allocate pages beyond the current pointer, the allocation must stall until the current pointer moves right enough to cover the required pages. Thus, a key trade-off in this design is that: the number of duplicated pages—i.e., the region defined by end pointerstart pointer—between GPU and CPU memory. A smaller duplication window reduces the duplication ratio but increases the likelihood of allocation stalls. For instance, if only one page is duplicated and an inference query requires three pages, the query must wait until the current and end pointers advance by two additional pages. On the other hand, if three pages are duplicated, the same query can proceed immediately without stalling, though at the cost of higher duplication ratio.
5 Standardized Interface for Connecting the KV Caching Layer and Inference Engine
Modern LLM inference engines, such as vLLM and SGLang, evolve rapidly to support newly released models with diverse architectures. For example, in 2025, an average of 15–20 new models are released each week. Supporting these new architectures often requires non-trivial modifications to inference engines, such as adding support for Sliding Window Attention or Multi-Head Latent Attention. These code changes frequently alter how KV cache is managed internally, making it infeasible for LMCache to adapt in an ad-hoc manner.
To address this challenge, LMCache introduces a standardized KV cache connector interface that decouples KV cache management from the inference engine backend. This design ensures that LMCache remains compatible regardless of how the underlying inference engine evolves.
Concretely, in vLLM, LMCache integrates through two key interfaces: 1) the scheduler, where the number of prefill tokens directly influences scheduling decisions and are changed by LMCache (i.e., if there is cache hit in LMCache, the number of tokens that need to be newly prefilled changes); and 2) the model runner, where inference computation occurs, and where KV cache loading and storing must be performed before and after execution.
The remainder of this section lists all the interfaces in Table 2, discusses the design for important APIs, and then traces how a query interacts with these interfaces end-to-end.
The interfaces listed in Table 2 form the foundation of LMCache ’s KV cache loading and storage across lower-tier storage. Among them, the first three interfaces are implemented within the vLLM scheduler, where they prepare the necessary metadata based on the number of matched tokens found in LMCache ’s KV cache backend. The remaining four interfaces reside in the model runner, which is responsible for executing the actual KV cache transfers between the inference engine and LMCache ’s KV cache backend.
Putting it together, when a query comes in, the scheduler first calls get_num_new_matched_tokens which queries LMCache to see cache hit tokens in the backend. Then the update_state_after_alloc function decides whether each page in vLLM needs to be loaded from external storage backend based on matched tokens information from LMCache. If the cache hit tokens are greater than zero, build_connector_meta function is called to prepare necessary metadata to load or store KV cache from storage devices.
Once the query reaches the model runner, in the case of layerwise pipelining, start_load_kv is called to start loading KV cache of the first layer to GPU memory. Then before each layer’s LLM inference computation starts, wait_load_kv is called to synchronize the KV cache loading for this layer, and starts the KV cache loading for the next layer. After each layer’s inference computation, in the layerwise case, wait_store_kv is called to wait until the KV cache for the previous layer has finished storing, and then calls start_store_kv to start the storing of KV cache for the newly generated KV cache layer.
In the case of non-layerwise pipelining, before the first layer’s LLM inference starts, start_load_kv is called to load the entire KV cache to GPU memory in a blocking manner. LLM inference will happen after the KV cache is put to the right GPU memory paged addresses. Then after the LLM inference has done for the current scheduling iteration, start_store_kv is called store the generated KV cache to lower-tier storage synchronously.
Interface Description
lookup(tokens) $\rightarrow$ {instance_id: hit_tokens} Returns the instance names of KV caches containing prefix-match tokens with the given token list.
query_ip(instance_ids) $\rightarrow$ IP Returns the IP addresses for the corresponding instance_ids
move(source, destination, tokens) Moves the KV cache of the given token list from a source instance to destination.
clear(tokens, instance_id, storage_device) Clears the KV cache for corresponding tokens from the storage_device in instance.
pin(tokens, instance, storage_device) Pins the KV cache for corresponding tokens at the storage_device in instance.
compress(tokens, instance, storage_device, compression_method) Compresses the KV cache for the corresponding tokens with a compression method and store it at the storage_device in instance.
Table 3: Functions in LMCache’s connector.
6 Controller Interfaces
There are many scenarios where higher-level applications need to see KV cache locations and operate on that. For example, query routing can benefit from knowing where the largest number of prefix cache hit tokens is at. To support such use cases, LMCache exposes APIs, as shown in Table 3, for querying and manipulating KV cache, such as locating cached entries or migrating them across instances.
The KV cache controller in LMCache consists of two components: a centralized manager and per-instance workers. Incoming controller queries are handled by the centralized manager, which dispatches the appropriate operations to workers running within each LMCache instance. The remainder of this section illustrates how to use the controller interfaces through several example applications.
KV cache–aware routing:  In this case, higher-level query routers would like to direct queries to instances with the largest number of prefix cache hits. The router first calls lookup(tokens), where the centralized manager dispatches the lookup queries to individual LMCache instances, to identify the instances containing matching prefix tokens, then invokes query_ip(instance_ids) to map instance ids to IP addresses. Finally, the query is forwarded to the IP with the highest number of hits.
KV cache migration:  When an instance holding KV cache fails or load balancing is required, the KV cache may need to be migrated across instances. The move(source, destination, tokens) interface moves the KV cache corresponding to tokens from the source instance to the destination.
KV cache clearance:  Applications may clear cache when switching models or reclaiming memory. The clear(tokens, instance, location) function removes the KV cache corresponding to tokens stored at a specific storage location (e.g., GPU memory or CPU memory) within the given instance.
Pinning KV cache in GPU memory:  Some contexts, such as system prompts in chatbot applications, should remain in GPU memory for fast access. The pin(instance, location, tokens) function pins the specified KV cache in the chosen storage tier (GPU memory) of the given instance.
There are other more advanced interfaces users can call to manipulate the KV cache, such as the compress(tokens, instance, location, compression_method) function which compresses KV cache stored in a specific location in a LMCache instance with compression_method so that the KV cache takes smaller amount of space in storage. Users can freely call these interfaces to explicitly manage the KV cache in their own applications based on their needs.
7 Evaluation
Figure 5: Single-node evaluation results.
Figure 6: Real-trace evaluation results.
Figure 7: Remote backend evaluation results.
Figure 8: TTFT and ITL CDF comparison for PD.
Figure 9: Comparing to SGLang CPU Offloading result.
Figure 10: With request asynchronization, LMCache overlaps KV cache loading and inference computation (either prefill or decode).
Figure 11: Latency breakdown of PD.
7.1 Setup
Scenario Acronym
Single-node /
Multi-node
Network
Medium
Real-world
Examples
CPU Offload Single-node N.A.
Single-node
CPU Offloading
Central Storage Single-node Ethernet
Centralized
Storage Server
PD Single-node NVLink
PD
Disaggregation
Table 4: Evaluation scenarios setup.
We evaluate LMCache under three different scenarios, as shown in Table 4. The three scenarios are representative setups that are commonly used by the users of LMCache.
Models:  We compare LMCache against baseline solutions on popular open source models adopted by industry: meta-llama/Llama-3.1-8B-Instruct, meta-llama/Llama-3.1-70B-Instruct, Qwen/Qwen2.5-Coder-32B-Instruct, Qwen/Qwen3-Coder-480B-A35B-Instruct-FP8, Qwen/Qwen2.5-72B-Instruct.
Datasets:  LMCache is evaluated on several datasets, including emulated multi-round question answering, long context question answering from LongBench Bai et al. (2024), and random dataset from vLLM official benchmarking script Kwon et al. (2023c).
[…]
Metrics:  For each experiment, we show both time-to-first-token (TTFT), which is the prefill delay, and inter-token-latency (ITL), which is the average delay between the generation of two consecutive output tokens. For component-wise analysis which breaks down the delay for CPU offloading or PD disaggregation, we report the delay for each component separately.
Baselines:  We compare LMCache with several baselines, including 1) vLLM with prefix caching, which implements prefix caching, but only keeps limited amount of KV cache inside GPU memory; 2) commercial offering A, B, and C, which provide dedicated endpoint service that reserves GPUs for users to run a user-defined model.
7.2 Single-node CPU Offloading
We first evaluate LMCache on the CPU Offload scenario as in Table 4. In this experiment, we use multi-round Q&A workloads that emulate a typical chatbot-based document analysis scenario. By default, each LLM query contains 10K tokens, consisting of a document (roughly a 12-page PDF) used as context and a unique short question. Llama-3.1-8B-Instruct model takes 20K tokens as input, since smaller models are generally better can handle more and longer queries. The LLM output is a short answer of 100 tokens at max. The chat session begins with 40 users, and additional users join according to a specified arrival rate (QPS). We set the maximum CPU memory LMCache can offload KV cache to 500 GB.
[…]
Understanding LMCache’s gains:  LMCache outperforms baselines for several reasons. Compared with naïve vLLM prefill, which caches KV data only in GPU memory, LMCache leverages CPU offloading. Since CPU memory can hold far more KV cache than GPU memory, LMCache achieves significantly higher cache hit ratios. With its efficient CPU-to-GPU loading scheme, cached data can be fetched quickly to accelerate inference. Our comparison with closed-source commercial alternatives is conducted in a black-box manner since their internal implementations are not publicly available. From the end-to-end results, we hypothesize that Commercial Option #1 lacks a KV cache offloading mechanism to secondary storage. In contrast, Commercial Option #2 likely supports KV cache offloading to secondary storage, yet its performance is still worse than LMCache.
7.3 Centralized Storage Server
Next, we run LMCache for KV cache sharing through a centralized remote server, that is connected to the GPU instance with a bandwidth of 15 Gbps, following the setup of central storage in Table 4. For this experiment, we evaluate using the TriviaQA dataset from LongBench Bai et al. (2024), a widely adopted benchmark for long-context evaluation. We follow the official vLLM benchmarking scripts Kwon et al. (2023c), which generate inference queries according to a Poisson distribution at a specified QPS.
[…]
We note, however, that loading KV cache from the remote backend introduces greater latency than loading from CPU memory, since the remote backend has a much lower bandwidth. As a result, the loading delay may even surpass the prefill delay, particularly when the input context is short or model is small, as prefilling is too fast in such cases—a scenario that we will demonstrate later in 7.7.
7.4 PD disaggregation
[…]
The performance gains of LMCache over the baseline stem from its more efficient design for PD disaggregation. Specifically, LMCache copies each chunk of the KV cache (generated during chunked prefill) to a buffer in the GPU memory of the prefiller instance, and then transfers it to the corresponding buffer on the decoder instance. Once received, the KV cache is copied into the paged memory of the decoder instance.
In contrast, vLLM’s native PD disaggregation sends the paged KV cache generated by the prefiller directly to the decoder, using NIXL’s memory copy function. This function takes as input the memory addresses of the KV cache pages on the prefiller side and copies them to the destination addresses on the decoder side. However, when the paged memory for the KV cache is scattered across the prefiller’s GPU memory, the transfer is performed in a page by page manner, which leads to bandwidth underutilization, as discussed in §4.
7.5 Real-trace Driven Evaluation
We also evaluate LMCache on a real production trace provided by company FF [^4] Anonymized due to NDA.. The trace includes query arrival timestamps and actual multi-round chat inputs. Notably, later rounds within the same user session often reuse conversation history from earlier rounds, resulting in prefix reuse. On average, the inputs in the trace are about 4K tokens long.
Since we do not have access to company $F$’s proprietary models, we replay the trace using the Sao10K/L3-8B-Lunaris-v1 model. To make the experiment tractable, we stretch the original trace which lasts for several days and apply random sampling so that the workload we run completes within one hour.
As shown in Figure 6, LMCache consistently outperforms naive vLLM on the real trace on different QPS, reducing both TTFT and ITL, demonstrating its effectiveness in real-world settings. Note that for lower QPS from 2 to 5, LMCache still outperforms naive vLLM by 25%, while when QPS is 6, LMCache has improvement of 49% compared to naive vLLM.
7.6 Component-wise Evaluation
To further understand the gain brought by LMCache, we also perform component-wise analysis to break down the delay of each component in the end-to-end system.
PD disaggregation:  Figure 11 shows the latency breakdown of LLM inference, including both prefill and decode computation, as well as the transmission of KV cache between prefiller and decoder instances. The prefill and decode computation times are the same for LMCache and vLLM’s native PD disaggregation. However, as discussed in §7.4, vLLM’s native design transmits KV cache at a much finer granularity, which results in bandwidth underutilization. In contrast, LMCache employs a more efficient KV cache transfer mechanism, enabling significantly faster transmission and thereby reducing the overall end-to-end delay in PD disaggregation.
Method Achieved Bandwidth
LMCache 400 Gbps
vLLM’s Native CPU Offloading 88 Gbps
Table 5: LMCache achieves much higher loading bandwidth when loading KV cache from CPU memory, compared to vLLM’s native CPU offloading.
CPU offloading:  In Table 5, we perform an ablation study to test the achieved loading bandwidth from CPU for LMCache and vLLM’s native CPU offloading. The reason LMCache achieves higher transfer bandwidth than vLLM’s native CPU offloading is due to the transfer granularity. While native CPU offloading performs data movement page by page, LMCache transfers data chunk by chunk. Each transfer operation triggers a CUDA memory copy, which involves preparing metadata beforehand and sending a completion signal afterward. These per-transfer operations add overhead to every memory copy kernel. By transferring larger chunks of data per copy, LMCache reduces the overall overhead, resulting in a much higher effective bandwidth.
[…]
7.7 Sensitivity Study
We also perform several sensitivity evaluation to see how LMCache’s delay changes under different context lengths and different types of remote backends.
Figure 12: Impact of different context lengths.
Impact of context lengths:  Figure 12 shows the prefill delay on B200 machines and the loading delay under different network bandwidths. When the network bandwidth is low (i.e., 32 Gbps), LMCache’s KV cache loading outperforms naive prefilling only when the input context length exceeds 256K tokens. In contrast, when the bandwidth is higher (i.e., 64 or 128 Gbps), LMCache’s loading consistently achieves lower delay than naive prefilling across all context lengths. These results suggest that LMCache’s KV cache loading should be adaptive: under low bandwidth, loading should be enabled only when the context length surpasses the crossover point where loading becomes faster than prefilling.
7.8 SGLang Results
Although our primary evaluation uses vLLM, we also evaluate LMCache integrated with SGLang. Figure 9 reports results for Qwen3-32B served on two H100 GPUs (TP=2) with LMCache ’s CPU offloading enabled. Compared to SGLang without CPU offloading, LMCache achieves higher throughput and lower mean TTFT and mean end-to-end latency. Compared to SGLang’s native CPU offloading, LMCache achieves comparable performance. These results confirm that LMCache is also effective on another inference engine. Although SGLang’s native CPU offloading achieves performance comparable to LMCache ’s CPU offloading on SGLang, it lacks a distributed storage backend capable of efficiently offloading data across a hierarchical set of storage devices, such as local disks and remote CPU/disk resources.
8 Real-World Lessons and Experience
8.1 Scaling KV Cache Storage and Offloading
One clear trend in the adoption of LMCache in production systems is that more and more KV cache needs to be stored, definitely beyond the number that can be stored in GPU memory. Early deployments kept all KV caches in GPU memory, but long contexts and multi-user workloads may quickly exhaust GPU memory. Thus, many companies now need to offload KV data to CPU RAM, CPU memory pooling, or even disks when needed. This prevents evicting useful cache entries — without offloading, a follow-up query would trigger full prefill of the same input context, often long contexts, which greatly prolongs TTFT.
Interestingly, some teams, including company R, have experimented with remote storage devices. A remote data fetch is orders of magnitude slower than GPU memory access, yet real-world tests have shown it can still yield both speedups and cost savings. Retrieving previously computed KV cache over the network can be faster than recomputing them on loaded GPUs, with LMCache’s KV cache loading optimizations (§4). While the overhead of loading KV cache from remote is large, it can be pipelined with inference computations, either prefilling requests that do not have cache-hit tokens or decoding computations. These experiences show that even slow and cheap storage can be leveraged to efficiently store and load KV caches.
Another technique used in production is the combination of CPU offloading and PD disaggregation. At company T, on the prefiller side, the KV cache is not only sent to the decoder side, but is also offloaded to the CPU memory of prefiller instance to achieve both the benefit of CPU offloading and PD disaggregation.
8.2 Emerging Use Cases and Surprising Outcomes
While chatbots were one of the first obvious beneficial usage scenarios of KV cache reuse, new use cases are rapidly emerging. Recently, large-scale recommendation systems from company L and company M have also adopted LLMs as the recommender model and becomes an important use case of KV cache reuse. In these systems, LLMs are used as the embedding models to prefill the contexts, without generating tokens, and are frequently reused by different queries of the same user. Furthermore, the contexts are often very long. Caching those KV cache can directly eliminate the expensive prefill phase, thus is a perfect use case of LMCache.
Although multi-round chat has long been a canonical use case for KV cache reuse, real-world deployments have revealed several important insights. Many systems initially adopted sliding context windows to fit more tokens into limited GPU memory during inference. However, this approach often degrades generation quality by discarding useful history and reduces prefix cache hit ratio, since truncated contexts are no longer prefixes. Deployment experiences at Company T and F demonstrate that keeping the full conversation history, while offloading KV caches to larger memory tiers, is a more effective practice. By avoiding truncation, the system achieves significantly higher cache hit rates on shared prefixes, leading to less GPU memory usage per request. Moreover, in production environments, companies have observed that prefix cache hit rate is far higher than they expected.
The experience from Company B’s adoption shows that industry users are accepting lossy optimizations, such as KV cache compression, for better system performance. Especially, in open-ended chatbot applications, where there is no single “correct” answer, using compressed KV cache (e.g., quantized KV cache) may not harm the user experience but can dramatically reduce memory usage and I/O time. Even in finance companies, teams found that slight differences from using a compressed KV cache are tolerable if the overall system throughput improves.
8.3 Integration and Deployment Challenges
Deploying a caching layer in production LLM inference systems has its own practical challenges and lessons. One key lesson we found in the adoption of LMCache is that many companies preferred containerized environments, such as Docker images, when deploying the caching layer into production, as they often do not go into the source code of LMCache.
Another important aspect in the caching layer for LLM inference is fault tolerance and transparency. Because the caching layer sits beneath the inference engines and is supposed to be invisible to end users, robustness is important. Many companies stressed that if the cache system fails or underperforms, it should not bring down the model service. For instance, LMCache incorporates fault-tolerant KV cache retrieval: if errors occur during retrieval, LMCache reports which portions of the KV cache were successfully read, ensuring the inference engine does not break due to the unsuccessful read of KV cache and the correctness of generation is preserved.
We also observed a growing demand for a clean separation between the model serving engine and the caching engine. Many storage companies, such as company I, R, C, and W, have explicitly asked for decoupling the KV cache management and the core inference code, since they want to touch as minimum core inference logic as possible. Although LMCache does not solve this for now, the trend is clear: an LLM inference engine is pairing with a variety of external KV cache management services.
At last, we found that ensuring compatibility between LMCache and emerging inference frameworks was easier than expected. LMCache was originally designed and developed on top of vLLM. However, integrating LMCache with other frameworks, such as SGLang, has also been straightforward, as long as they adopt the paged attention mechanism. Looking ahead, we expect that LMCache will be equally easy to integrate with other inference frameworks based on paged attention.
8.4 Adoption, Lessons Learned, and Future Outlook
Our journey began as an academic prototype in the summer of 2024. At that time, KV caching was a niche concept mostly discussed in a few research papers. But by spring 2025, the landscape had changed — KV cache became a key technique in industry for improving the efficiency in LLM inference. The adoption of LMCache truly exploded in early 2025, with numerous companies starting to collaborate and use.
Interestingly, although LMCache has initially been an academic prototype, we found that LMCache has much more industry users than academia ones. One potential reason is that there is lack of an easy set of programmable interfaces. As a concrete example, a rich set of research literatures explore dropping tokens from KV cache for compressing it into smaller tensors. However, it is non-trivial to implement token dropping in LMCache, since LMCache assumes the number of tokens that is input to and output of LMCache should be the same. Furthermore, any intermediate states in the attention module are hidden from the LMCache’s side, so research works that need to access or modify the attention states are hard to be implemented with LMCache.
Finally, we learned that programming languages for high performance are not always the best fit for building LLM inference infrastructure. Early on, some organizations (e.g.,, company R and I) assumed that KV cache management must be implemented in Rust or C++ to achieve high efficiency. In contrast, we developed LMCache in Python, which turned out to be a more effective choice. Python not only simplifies integration with existing inference frameworks but also lowers the barrier for community contributions, enabling rapid feature development and iterations. That said, CUDA-based implementations are still essential for certain high-performance data loading modules within LMCache.
9 Conclusion and Outlook
This paper presented LMCache, the first open-source and most widely adopted production-ready KV caching layer for enterprise-scale LLM inference. By treating KV cache as a first-class data structure rather than an internal byproduct of inference, LMCache transforms LLM engines from isolated token processors into a distributed ecosystem of compute and storage. Evaluation across diverse workloads and models demonstrates that LMCache consistently delivers significant throughput improvement and latency reduction compared to both open-source baselines and commercial inference APIs. Beyond performance, LMCache has already seen rapid adoption in production environments, where enterprises leverage its CPU offloading, hierarchical storage, and PD disaggregation capabilities to keep low latency and reduce cost in trillion-token–scale deployments. Real-world deployments have also revealed new opportunities, such as KV cache reuse in recommendation systems and lossy compression in open-ended chatbots, underscoring the versatility of LMCache across application domains.
Looking ahead, LMCache points to a broader shift: AI-native data such as KV caches will increasingly serve as the substrate for scaling LLM inference and agentic workloads. By establishing KV cache as a standardized storage and communication medium, LMCache lays the foundation for future systems that treat inference not as isolated sessions but as a persistent, cache-aware computation fabric. We hope that the design, optimizations, and deployment lessons presented in this paper will inform the next generation of LLM infrastructure, where AI-native data, such as KV caches, is not merely an optimization but a core primitive for efficient, reliable, and scalable inference.
The source code of LMCache is at: https://github.com/LMCache/LMCache.
10 Acknowledgement
We would like to thank the LMCache community for their invaluable support and contributions, including Baolong Mao and Chunxiao Zheng for managing remote connectors, Martin Hickey for GitHub Infrastructure, Huaizheng Zhang, Siddhant Ray, Zhuohan Gu and Hanchen Li for writing and maintaining documentation, Rui Zhang for the deployment of LMCache, Qizheng Zhang and Hussain Mohammad for insightful feedback.

### v2

@@
-
-Abstract
+LMCache: An Efficient KV Cache Layer for Enterprise-Scale LLM Inference
+LMCache: An Efficient KV Cache Layer for Enterprise-Scale LLM Inference
+Yuhan Liu12* Jiayi Yao12* Yihua Cheng1* Yuwei An1 Xiaokun Chen1 Shaoting Feng12
+Yuyang Huang12 Samuel Shen1 Rui Zhang1 Kuntai Du1 Junchen Jiang1
+1Tensormesh Inc. 2University of Chicago
+footnotetext: *Equal contribution
 […]
-footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
-*Equal contribution
 1 Introduction
 Today, large-language model (LLM) inference has outpaced training in growth.
-LLM inference powers a wide range of applications, from interactive customer support and code generation to retrieval-based document analysis and agentic workflows.
-However, the LLM inference systems have not caught up—latency (both response latency and generation speed) and cost are now the bottleneck.
-A key limitation of today’s LLM inference systems is that each user query is independently processed by one instance of the inference engine, without any data being reused across queries or inference engines.
-This means that a LLM query’s lifecycle, including computation and I/O operations, happens in the GPUs and the GPU memory of one inference engine.
-Each query takes a token sequence as input and another token sequence as output, and once a query completes, all outputs or intermediate states generated during inference are discarded.
-Figure 1: LMCache supports both context caching (KV cache offloading and sharing across queries) and PD disaggregation (cross-engine transfer of KV caches.
-This design, albeit simple, leads to redundant computation and resource underutilization, and two notable industry proposals attempted to address them (Figure 1).
-Cross-query caching to avoid redundant compute:  As each query is treated in isolation, the LLM must prefill the same context (prefix) each time the context is used, and the repeated prefill compute is expensive and causes slow response, as users cannot see the first token before the prefill completes.
-Yet, such redundant computation could be avoided if we persist KV cache of the prefix—which LLM created internally after it prefills the prefix—beyond the lifecycle of a query and reuse the KV cache when subsequent queries reuse the prefix (e.g., the same prompt template or retrieved document chunk).
-By avoiding redundant computation, cross-query context (prefix) caching significantly lowers both time-to-first-token (TTFT) and overall GPU resource consumption during the prefill phase.
-Prefill–decode disaggregation for higher utilization:  Co-locating the prefill phase (compute-bound, throughput-oriented) and the decode phase (memory-bound, latency-sensitive) in the same inference engine causes resource underutilization as the GPUs must be overprovisioned to ensure a consistently low decoding latency without being interrupted by prefill.
-In response, the industry attempted to decouple the throughput-oriented prefill processing from latency-sensitive decoding Such prefill-decode (PD) diaggregation architecture batches all queries’ prefill phase on one set of inference nodes and transfers their KV caches to another set of inference nodes that run the decoding phase only, thus mitigating tail decoding latency under high concurrency.
+LLM inference powers millions of applications, from interactive customer support and code generation to retrieval-based document analysis and agentic workflows.
+To build the systems for LLM inference, KV cache, the intermediate states of large language model (LLM) inference, has now become a de facto optimization to make inference faster.
+Traditionally, KV cache has been used during new token generation to skip re-computing the KV cache of the input prompts, within a single query.
+Thus, every query is processed independently by one instance of the inference engine, and a LLM query’s lifecycle, including computation and I/O operations, happens in the GPUs and the GPU memory of one inference engine.
+Figure 1: Upper: LMCache is being used by more and more users over time.
+Middle: LMCache is being used to store KV cache of larger size.
+Bottom: LMCache’s docker image pull counts have continued to increase.
+However, recent trends has advanced to propose moving KV cache outside the GPU memory, including two emerging directions:
+Cross-query caching to avoid redundant compute:  KV cache can be persisted (e.g., to lower-tier storage devices) beyond the lifecycle of a query to avoid any re-computation of the shared prefix for another query.
+Prefill–decode disaggregation for higher utilization:  There is an arising trend to decouple the prefill and decode phase on different GPUs, to ensure that latency-sensitive decoding phase is not affected by throughput-oriented prefill phase.
+Such prefill-decode (PD) diaggregation requires KV cache produced by the prefill GPUs to be transferred to the decode GPUs.
+Our real-world usage statistics confirm the trend of moving KV cache out from the GPU memory.
+As will be discussed in detail soon (§2.2), over time, the total size of KV cache that is stored by the users continues to grow, and it far exceeds the capacity of the GPU memory.
+This suggests that KV cache may need to be frequently evicted from GPU memory as more and more requests come in.
+Also, in order for it to be reused by another query, it needs to be offloaded from GPU memory, and then loaded back to GPU memory for reuse.
 To make these ideas practical, the LLM inference systems must be augmented with new KV cache semantics.
@@
 The system also must allow the extracted KV caches to be stored persistently and transferred across distributed inference engines.
-Most importantly, for such interface extensions to be practical, KV cache extraction, re-loading, storage, and transfer must be efficient, and the new interfaces must remain compatible with rapidly evolving inference engines such as vLLM Kwon et al.
-(2023b) and SGLang Zheng et al.
-(2024).
+Most importantly, for such interface extensions to be practical, KV cache extraction, re-loading, storage, and transfer must be efficient, and the new interfaces must remain compatible with rapidly evolving inference engines such as vLLM vllm and SGLang zheng2024sglangefficientexecutionstructured.
 We introduce LMCache, the first open-source library that provides a high-performance implementation of these new KV cache semantics.
@@
 […]
-The remainder of this paper details the motivation (§2), LMCache architecture and key design choices (§3–§6), experimental evaluation (§8), and deployment experiences (§9).
-2 Motivation
-2.1 Inference Cost and Delay as the Bottleneck
-The past two years have witnessed explosive growth in the deployment of LLMs across diverse enterprise applications.
-As the user base and query volume of LLM-based applications increase, the cost and delay of inference, rather than training, emerges as the primary bottleneck to scaling LLM serving to production scale.
-Inference cost:  Besides the simple fact that LLMs are queried by more users, the increasing input length and output length also contribute to raising inference cost, as each input token and each output token incur an additional cost to process.
-Four trends are behind the increase in input and output lengths.
-(i) More interactions by each user with LLMs build a longer user history, which will be prepended to future LLM inputs as the context.
-(ii) LLMs increasingly take multimodal inputs, such as images and videos, which typically will be transformed to a long sequence of tokens before being fed to LLMs.
-(iii) Newer LLMs often have longer context windows, allowing context/prompt engineers to simply stuff more tokens into LLM inputs.
-(iv) Finally, reasoning models produce unusually long output, which might be used as input to subsequent queries in one workflow.
-Even with steady improvement on GPU price and performance, serving a trillion-token-scale workload can incur millions of dollars in annual compute expenditure Nie et al.
-(2024); Databricks Research (2024).
-Redundant computation across queries (e.g., recomputing the same prompt prefix hundreds of times) exacerbates these costs Zhang et al.
-(2024); Caylent (2024).
-Inference delay:  In interactive applications, tail latency (e.g., 95th/99th percentile) is as critical as median latency.
-Slow responses in the worst case, often due to resource contention or large context processing, directly degrade user experience and can lead to user abandonment, in much the same way video streaming stalls affect viewer experience Liu et al.
-(2024a).
-In particular, empirical observations from industry reports and academic studies indicate that:
-  • $\bullet$
-    Time-to-first-token (TTFT) can be dominated by the prefill phase for long-context inputs, especially when context lengths exceed 8–16k tokens.
-  • $\bullet$
-    Inter-token latency (ITL) remains low for single-stream decoders, but rises sharply when GPU compute is shared across many concurrent sessions.
-  • $\bullet$
-    Tail delays are disproportionately impacted by scenarios requiring large prompt recomputation or reallocation of memory-bound resources.
-2.2 KV Caching as a Cross-Cutting Optimization
+Beyond quantitative gains, LMCache has seen adoption across several enterprises and open-source projects, providing useful insights and lessons in KV cache-driver optimizations at production scale.
+This includes surprising gain in latency brought by remote storage backend, reduction in prefix cache hit rate because of context truncation, and that the flexibility to evolve faster and integration smoothness is more important than the language performance itself.
+The remainder of this paper details the motivation (§2) and challenges (§3), LMCache architecture and key design choices (§4, §5, and §6), deployment experiences (§9) and experimental evaluation (§8).
+Figure 2: LMCache supports both context caching (KV cache offloading and sharing across queries) and PD disaggregation (cross-engine transfer of KV caches.
+2 Motivation and Real-world Usage Statistics
+2.1 KV Cache in LLM Inference
 KV cache was originally introduced to accelerate a single inference query by storing the attention states, in the form of $K$ and $V$ tensors, for input tokens and previously generated tokens directly in GPU memory.
@@
 In short, it is a LLM-native representation of knowledge.
-In all popular transformer architectures, these caches are stored and used only within the lifecycle of an individual query and discarded when the query ends.
-Yet, KV caching can be extended to support two critical performance optimizations in production-scale serving:
+Nowadays, the contexts have grow longer and longer, and people have started to augment inference with background knowledge.
+Given this trend, it is popular to share KV cache across different user queries to reduce the redundant computations for the long contexts or background knowledge.
+Figure 3: Weekly growth of KV cache size, including portions that fit in GPU memory and those that exceed it.
+2.2 Real-world Usage Statistics
+KV Cache Size Exceeds GPU memory:  Although KV cache has been kept inside GPU memory for all traditional LLM inference systems, we observe that the required size of KV cache is now far exceeding the GPU memory capacity, from our real-world usage statistics, gathered by usage tracker voluntarily turned on by users.
+Figure 3 shows the weekly growth of KV cache size over the past five weeks, for caches that fit within GPU memory (green) and those that exceed GPU memory capacity (blue).
+The portion of KV cache that no longer fits in GPU memory has increased significantly over time, showing that GPU memory alone is insufficient for storing all caches.
+To enable KV cache reuse across queries, especially those generated long before reuse, it becomes necessary to move KV cache out of GPU memory, for instance, by offloading it to CPU memory or other storage tiers.
+Reuse per Token has Greatly Increased:  We also observe that reuses per token has greatly increased over time.
+As shown on the left side of Figure 4, where we plot the ratio between reused tokens and all stored tokens, beyond the GPU memory, plotted with top-10 users.
+We denote this ratio as reuses per token.
+The reuses per token has grown significantly over the past several weeks, which indicates that tokens that cannot fit inside GPU memory are being more and more frequently reused by inference.
+This suggests that more and more tokens need to be loaded back to GPU memory.
+On the right hand side, the figure shows the distribution of reuse per token for different users over the past week.
+More than 19% of users reuse stored tokens for more than 1.5 times, suggesting the trend of users accessing a token multiple times after it is stored.
+Figure 4: Left: Average reuse per token for top users. Right: Distributions of average reuse per token across different users.
+2.3 Need an Efficient KV Caching Layer for Moving KV Cache
+From the above observations from the statistics gathered in real-world deployments, we find two important trends of KV caching.
+First, the KV cache that cannot simply fits in GPU memory keeps growing, potentially due to growing length of the contexts or larger amount of user traffic.
+Second, reuses per token stored beyond the GPU memory has also increased over time.
+Both trends suggest that we need to move KV cache out of GPU memory.
+Specifically, in the current industry, two scenarios which move KV cache out from GPU exist:
   • 1.
@@
 Examples include document analysis where the same document (chunk) remains constant across multiple queries, and multi-turn dialogues with a fixed system prompt or long preamble.
-Prefix caching reduces redundant computation during the prefill phase, directly lowering TTFT and GPU-hours per query Liu et al.
-(2024b); Gao et al.
-(2024); Jin et al.
-(2025b); Ren et al.
-(2025); Qin et al.
-(2025a); Jin et al.
-(2024); Chen et al.
-(2024; 2025).
+Prefix caching reduces redundant computation during the prefill phase, directly lowering TTFT and GPU-hours per query liu2024cachegenkvcachecompression; gao2024costefficientlargelanguagemodel; cake; 10.1145/3719330.3721230; mooncake; jin2024ragcacheefficientknowledgecaching; chen2024kvdirectdistributeddisaggregatedllm; impress.
   • 2.
     Prefill–decode (PD) disaggregation (i.e.,cross-engine KV cache transfer) splits inference into a prefill stage (processing the entire input prompt) and a decode stage (autoregrssive token generation) across different GPUs or nodes.
-This approach reduces tail latency by maximizing the decoding speed without being interrupted by the prefill phase Zhong et al.
-(2024); Patel et al.
-(2024); Shi et al.
-(2025).
-Roles of KV caching:  Both optimizations critically depend on the ability to move KV cache segments efficiently to and from GPU memory, CPU DRAM, NVMe storage, or over the network.
-  • $\bullet$
-    Context reusing stores the KV cache of a context for as long as the context will be reused, so it requires a hierarchical storage of KV caches and efficient sharing of KV caches across multiple serving engine instances.
-  • $\bullet$
-    PD disaggregation requires fast KV cache transfer between GPUs, often across PCIe, NVLink, or RDMA, in order to minimize the end-to-end delay of each query, which runs across the prefiller nodes and the decoder nodes.
+This approach reduces tail latency by maximizing the decoding speed without being interrupted by the prefill phase distserve; splitwise; shi2025nexusproactiveintragpudisaggregationprefill.
+However, there lacks a library to support efficient extraction and loading from and to the GPU memory due to the system challenges as discussed soon (3).
+3 Challenges of Efficient KV Caching and Related Work
 3.1 Challenges of Efficient KV Caching
@@
 100MB 49GBps
-Table 1: Transfer message size vs achieved transfer throughput using RCCL transfer library UCCL Team (2025).
+Table 1: Transfer message size vs achieved transfer throughput using RCCL transfer library UCCL2025KVTransferEngine.
 3.1.1 Challenge #1: I/O inefficiency under paged memory
@@
 These methods introduce non-trivial delay overhead, especially when handling large data structures like KV caches, and lack zero-copy support with various storage devices (local or remote), causing extra CPU-GPU data copies.
-Recent high-throughput inference engines, such as vLLM Kwon et al.
-(2023b) and SGLang Zheng et al.
-(2024), make KV cache storage and transfer even more challenging.
+Recent high-throughput inference engines, such as vLLM vllm and SGLang zheng2024sglangefficientexecutionstructured, make KV cache storage and transfer even more challenging.
 They employ paged attention memory, dividing the attention buffer into small, fixed-size pages (typically 16–64 KB).
@@
 However, because the pages of a KV cache are not always contiguous, the paged memory architecture dramatically increases the number of small-sized I/O operations required to persist or transfer a KV cache.
-Transferring such small chunks of data is known to suffer from network bandwidth underutilization and reduce throughput Kwon et al.
-(2025); NVIDIA Developer Forums (2020); Meta Engineering (2024).
-Prior work (Table 1) has shown that, on a setup with two AMD GPU nodes connected by eight Broadcom Thor-2 400Gbps NICs, the transfer size must reach at least 16 MB to saturate the available network bandwidth Zhou et al.
-(2025).
-Furthermore, prior work has shown that only transferring a data size of megabyte range (e.g., 1–2MB) can achieve 75–80% of the theoretical PCIe 5.0 bandwidth Xie et al.
-(2025).
+Transferring such small chunks of data is known to suffer from network bandwidth underutilization and reduce throughput kwon2023demystifyingnccl; nvidia2020smalltransfer; facebook2024roce.
+Prior work (Table 1) has shown that, on a setup with two AMD GPU nodes connected by eight Broadcom Thor-2 400Gbps NICs, the transfer size must reach at least 16 MB to saturate the available network bandwidth uccl_transport.
+Furthermore, prior work has shown that only transferring a data size of megabyte range (e.g., 1–2MB) can achieve 75–80% of the theoretical PCIe 5.0 bandwidth strata.
 3.1.2 Challenge #2: Compatible with fast-evolving inference engines
 With the widespread use of AI, new LLMs and hardware accelerators are introduced at a rapid pace.
-In 2025, one prominent LLM was released on average every 4 days bes (2025).
+In 2025, one prominent LLM was released on average every 4 days best44llms2025.
 In response, inference engines must evolve just as quickly.
@@
 Several KV cache handling mechanisms exist, but none of them fully address the above challenges:
-Inference frameworks:  Since the release of vLLM Production Stack vLLM project (2025) in January 2025, there have been several open-source distributed inference stacks, including Nvidia’s Dynamo NVIDIA Corporation (2025), AIBrix Team et al.
-(2025), llm-d llm-d Project (2025), SGLang OME Team (2025), and KServe Contributors (2025b).
+Inference frameworks:  Since the release of vLLM Production Stack vllm-production-stack2025 in January 2025, there have been several open-source distributed inference stacks, including Nvidia’s Dynamo aiDynamoDynamo2025, AIBrix theaibrixteam2025aibrixscalablecosteffectivelarge, llm-d llmDllmD2025, SGLang OME ome, and KServe KServe2025.
 They focus on easy deployment of inference engine solutions over Kubernetes, and technically, they all support various query routers based on load or prefix cache awareness and support KV caching, where LMCache is used in vLLM production stack, Dynamo, llm-d, and KServe.
@@
 We will evaluate their performance and compare it against LMCache in §8.
-KV cache storage layers:  Mooncake Qin et al.
-(2025b), Redis Redis (2025), InfiniStore ByteDance (2025), and 3FS Contributors (2025a) provide distributed object storage or caching, but they lack an efficient “glue” layer between the inference engines to efficiently move small tensors frequently across different storage tiers, or are tightly coupled with a specific inference framework.
+KV cache storage layers:  Mooncake qin2025mooncakekvcachecentricdisaggregatedarchitecture, Redis redis-enterprise-software-references, InfiniStore bytedanceInfiniStore2025, and 3FS DeepSeek2025ThreeFS provide distributed object storage or caching, but they lack an efficient “glue” layer between the inference engines to efficiently move small tensors frequently across different storage tiers, or are tightly coupled with a specific inference framework.
 Proprietary implementations:  Proprietary inference APIs (e.g., Fireworks AI, Together AI) implement their own prefix caching internally, but these are tied to their closed-source serving stacks and are not accessible to operators deploying their own infrastructure.
-Source code for research:  Several research proposals have open-sourced prototypes for their KV cache optimizations, including prefix caching Kwon et al.
-(2023a); Zheng et al.
-(2024); Yu et al.
-(2025); Gim et al.
-(2024); Ye et al.
-(2024); Lee et al.
-(2024); Zhao et al.
-(2024); Jin et al.
-(2024); Gao et al.
-(2024); Chen et al.
-(2025); Jin et al.
-(2025a); Yang et al.
-(2025), PD disaggregation Zhong et al.
-(2024); Patel et al.
-(2024); Shi et al.
-(2025), and KV cache compression Liu et al.
-(2024c); Jegou et al.
-(2024); Xiao et al.
-(2024a; b); Li et al.
-(2024); Qin et al.
-(2025c); Tang et al.
-(2024); Ge et al.
-(2024); Li et al.
-(2025); Du et al.
-(2025); Zhang et al.
-(2025).
+Source code for research:  Several research proposals have open-sourced prototypes for their KV cache optimizations, including prefix caching vllm; zheng2024sglangefficientexecutionstructured; pensieve; prompt-cache; chunkattention; infinigen; blendserve; jin2024ragcacheefficientknowledgecaching; gao2024costefficientlargelanguagemodel; impress; cake; kvshare, PD disaggregation distserve; splitwise; shi2025nexusproactiveintragpudisaggregationprefill, and KV cache compression liu2024kivi; kvpress2024; xiao2024duoattentionefficientlongcontextllm; xiao2024efficientstreaminglanguagemodels; snapkv; qin2025cakecascadingadaptivekv; quest; ge2024modeltellsdiscardadaptive; commvq; bitdecoding; pqcache.
 However, these prototypes are typically built on research-oriented inference framworks, such as HuggingFace Transformers, not fully enterprise-ready, or are not designed to evolve alongside the rapidly changing inference engine ecosystem, such as SGLang and vLLM.
@@
 Figure 5: LMCache sits between LLM inference engines and heterogeneous storage/network devices.
-3.1 Architecture
-At a high level, LMCache connects to inference engines to move KV cache between GPU memory and the storage backend.
-Figure 3 shows the end-to-end system.
-Below, we walk through two example workflows: storing and retrieving KV cache.
+Figure 6 shows the end-to-end system. Below, we walk through two example workflows: storing and retrieving KV cache.
 Store:  When a new query arrives, it first passes through the KV connector, which prepares metadata such as the tokenized input prompt and GPU memory addresses of the relevant pages.
@@
 Figure 6: End-to-end system workflow for LMCache.
-3.2 Key Components
-LMCache is composed of a high-performance data plane and a flexible control plane.
-LMCache Worker (Data Plane).:  Each inference engine is paired with a LMCache worker.
-The LMCache worker module is responsible for moving KV caches between GPU memory and other tiers or other workers.
-It implements both KV offloading (to CPU or disk) and PD disaggregation (GPU–GPU transfer).
-To maximize throughput, workers employ kernel-optimized GPU buffers, asynchronous chunked I/O, and layer-wise pipelining, as elaborated in the next section.
-These techniques allow LMCache to sustain near-GPU-resident bandwidth even when handling small paged memory objects (16–64KB), a common configuration in vLLM and SGLang.
-LMCache Controller (Control Plane).:  The controller exposes a programmable API to system operators and higher-level schedulers.
-It provides cache lifecycle management primitives, such as pinning KV segments, compressing and decompressing cache tensors, migrating caches between devices, and evicting low-priority entries.
-The controller also maintains a virtualized namespace, enabling uniform addressing of caches across heterogeneous devices.
-Together, the LMCache worker and controller form a modular system that integrates seamlessly into enterprise-scale serving pipelines.
-The worker ensures fast data movement, while the controller provides higher-level control semantics necessary for efficient resource utilization.
 5 Performance Optimizations
@@
 , which are typically 20 KB–63 KB for popular models including Llama, Qwen, GPT-OSS etc.
-Such small units are inefficient for transferring, as they cannot saturate bandwidth Xie et al.
-(2025); Zhou et al.
-(2025);
+Such small units are inefficient for transferring, as they cannot saturate bandwidth strata; uccl_transport.
   • $\bullet$
@@
 .
-This is achieved using an intermediate GPU buffer.
-For storing, multiple pages (16 by default) are first copied from GPU memory into the buffer, then offloaded collectively to lower-tier storage (e.g., CPU memory) and saved at the granularity of chunks rather than individual pages.
-For loading, chunks are first retrieved from the storage layer into the GPU buffer and subsequently split into pages before being placed in the serving engine’s internal GPU memory.
-LMCache implements a customized CUDA kernel to accelerate the memory copy.
-Batched Store/load Operations:  LMCache supports storing and loading KV cache across multiple storage devices, including local CPU memory, local disks, and remote disks.
-In practice, KV caches often need to be transferred to different devices in parallel – for example, storing the KV cache of a hot context in CPU memory while simultaneously offloading a cold context to local disk.
-A naive approach that stores the two KV caches sequentially underutilizes available bandwidth, since transferring to CPU memory leaves the disk bandwidth idle.
-To address this, LMCache batches storage and loading queries across different tiers, thereby aggregating operations and fully utilizing the bandwidth between the GPU and multiple storage tiers.
-Delayed Decode KV Cache Storing:  LMCache also supports saving newly generated KV cache during decoding.
-In a naive design, each newly generated page is immediately offloaded to lower-tier storage, incurring frequent small writes.
-Instead, LMCache aggregates multiple pages and stores them together as a chunk, improving I/O efficiency.
+This is achieved using an intermediate streaming GPU buffer.
+For storing, the KV cache are first copied from the scattered paged GPU memory into a contiguous streaming buffer with a customized CUDA kernel, then offloaded collectively to lower-tier storage (e.g., CPU memory) with DMA engines at the granularity of chunks rather than individual pages.
+For loading, chunks are first retrieved from the storage layer into the GPU buffer with DMA engines and subsequently split into paged memory with CUDA kernels.
+Parallel Store/load Operations:  LMCache supports parallel storage and retrieval of KV caches across multiple storage tiers, including local CPU DRAM or disks, remote CPU DRAM or disks, and object storage (e.g., S3).
+In practical LLM serving workloads, KV caches often need to be migrated across devices concurrently—for instance, transferring the KV cache of a hot context from GPU to CPU memory while simultaneously offloading a cold context from CPU memory to local disk.
+To maximize link utilization, LMCache ’s store and load APIs accept multiple source and destination devices, enabling concurrent data movement across heterogeneous links.
+Moreover, these operations can be executed in parallel when the interconnect supports full-duplex communication (e.g., PCIe).
+Delayed Decode KV Cache Storing:  LMCache also supports storing newly generated KV caches during decoding.
+Instead of offloading each token’s KV cache immediately—a naive approach that triggers frequent small writes—LMCache buffers KV caches and performs batched storage once a predefined number of tokens (i.e., a chunk) have been generated.
+This chunk-based delayed storing strategy reduces write frequency, minimizes I/O overhead, and significantly improves overall storage throughput.
 5.2 Compute-I/O Overlapping
-LMCache’s data movement often happens in parallel to LLM inference computation.
-For example, during decoding, a KV cache is computed while being stored to storage devices at the same time.
-Naive data movement operations can block the LLM inference computations, or vice versa.
-LMCache employs multiple optimizations aiming for overlapping LLM inference computations with I/O.
+LMCache employs multiple optimizations aiming for overlapping LLM inference computations with I/O to maximize GPU utilization.
 Layer-wise pipelining:  LMCache overlaps KV cache transfers with inference computation through layer-wise pipelining.
@@
 Asynchronous compute & prefetch:  In many scenarios, there is a time gap between when the inference scheduler admits a query and when the query’s KV cache is actually needed for inference.
-For example, if 100 queries arrive simultaneously but the inference engine can only process 50 in parallel, the remaining 50 must wait in the queue before their inference begins.
-LMCache exploits this idle interval to prefetch the queued queries’ KV cache from slower storage tiers into faster ones (e.g., from remote disk to local CPU memory).
-As a result, when the actual inference computation starts, the required KV cache can be loaded directly from faster-tier storage, significantly reducing loading delay.
+For example, if a query (with cache hit) arrives when the inference engine is processing other queries, the arriving query has to wait in queue.
+LMCache exploits this idle interval to prefetch the queued queries’ KV cache from slower storage tiers into faster ones (e.g., from remote disk to local CPU memory or GPU memory).
+As a result, when the actual inference computation starts, the required KV cache can be loaded or used directly from faster-tier storage, significantly reducing loading delay.
 LMCache allows users to configure the target tier for prefetching based on their own needs in latency SLO and resource constraints.
-Process separation:  Co-locating LMCache ’s data movement with the inference engine in a single process introduces 5%–10% latency overhead due to CPU resource contention.
-To eliminate this overhead, LMCache decouples data movement into a separate process, isolating it from the inference computation process.
-Additionally, process separation also enables memory sharing across multiple inference engine instances.
-When data movement is co-located with computation, each inference instance maintains its own CPU memory pool, so instances cannot access one another’s KV cache without inter-process communications.
-By contrast, a standalone data movement process manages a unified CPU memory pool, allowing multiple inference instances to store and load KV cache from the same pool, thereby improving efficiency and reducing redundancy.
 5.3 Minimum Data Copy
@@
 LMCache avoids this by maintaining only the minimum required copies.
-Zero-Copy Operations:  When transferring KV cache to remote storage or another GPU instance (e.g., in disaggregated prefill), LMCache minimizes data duplication through a reference counter.
-Specifically, when KV cache is written to multiple destinations—such as local CPU memory, local disk, or remote disk—LMCache increments a reference counter for the shared data instead of creating new copies.
-Each completed write decrements the counter, and once the count reaches zero, the data is released.
-This design ensures that data is shared across concurrent write operations without unnecessary replication, thus reducing memory pressure and improving efficiency.
-This technique is simlar to PCB counter in operating systems Strecker (1978).
+Zero-Copy Operations:  When transferring KV cache to multiple devices simultaneously, LMCache minimizes data duplication through a reference counter.
+Specifically, when KV cache is written to multiple destinations—such as writing from local CPU memory to local disk and remote object storage at the same time—LMCache increments a reference counter on the shared data for each transfer instead of creating new copies.
+Each completed read or write decrements the counter, and once the count reaches zero, the data is released.
+This design ensures that data is shared across concurrent read and write operations without unnecessary replication, thus reducing memory pressure and improving efficiency.
+This technique is simlar to PCB counter in operating systems Strecker1978VAX.
 Figure 7: Illustration of dynamic offloading in LMCache.
@@
 Function name Description
-get_num_new_matched_tokens(query) $\rightarrow$ matched_tokens Returns the number of cache-hit tokens found in LMCache ’s backend.
+get_num_new_matched_tokens(query) $\rightarrow$ Optional[matched_tokens] Returns the number of cache-hit tokens found in LMCache ’s backend.
+Returns None if the LMCache decides to let vLLM process other requests first and put this request back to waiting queue.
 update_state_after_alloc(query, blocks, num_external_blocks) Updates whether a query needs to transfer KV cache from LMCache’s backend.
@@
 Note that if a query attempts to allocate pages beyond the current pointer, the allocation must stall until the current pointer moves right enough to cover the required pages.
-Thus, a key trade-off in this design is that: the number of duplicated pages—i.e., the region defined by end pointerstart pointer—between GPU and CPU memory.
+Thus, a key trade-off in this design is that: the number of duplicated pages—i.e.,, the region defined by end pointer – start pointer –between GPU and CPU memory.
 A smaller duplication window reduces the duplication ratio but increases the likelihood of allocation stalls.
@@
 On the other hand, if three pages are duplicated, the same query can proceed immediately without stalling, though at the cost of higher duplication ratio.
+Though not supported, the same dynamic offloading strategies can also be extended to other storage tiers beyong CPU and GPU.
 6 Standardized Interface for Connecting the KV Caching Layer and Inference Engine
@@
 To address this challenge, LMCache introduces a standardized KV cache connector interface that decouples KV cache management from the inference engine backend.
-This design ensures that LMCache remains compatible regardless of how the underlying inference engine evolves.
-Concretely, in vLLM, LMCache integrates through two key interfaces: 1) the scheduler, where the number of prefill tokens directly influences scheduling decisions and are changed by LMCache (i.e., if there is cache hit in LMCache, the number of tokens that need to be newly prefilled changes); and 2) the model runner, where inference computation occurs, and where KV cache loading and storing must be performed before and after execution.
+This design ensures that LMCache remains compatible regardless of how the upstream inference engine evolves.
+We note that the design of this API is initiated by LMCache team, but the implementation and the maintainance of this API are the collaborative effort of both LMCache team and vLLM team.
+Design objectives:  The key design objectives are:
+  • $\bullet$
+    Maximum flexibility: it enables as much KV cache operations as possible.
+  • $\bullet$
+    vLLM-native: it aligns with the design direction of vLLM, including strict scheduler — worker separation, prefix caching as the first-class citizen, and piece-wise CUDA graph, where vLLM only captures CUDA graph for non-attention operations.
+  • $\bullet$
+    Friendly to out-of-tree connector: it allow intergrating with out-of-tree connectors without vLLM-side code motification.
+  • $\bullet$
+    Minimum API-level overhead: it does not introduce overheads (e.g., inter-process communication) at API level.
+To be compatible with the scheduler — model runner separation philosophy in vLLM, the connector API contains two sets of interfaces: 1) the scheduler, where the extra cache-hit tokens from the connector are treated as normal prefix-cached tokens in vLLM and directly influences scheduling decisions and are changed by LMCache (i.e., if there is cache hit in LMCache, the number of tokens that need to be newly prefilled changes); and 2) the model runner, where we add hooks before and after model execution, and also before and after attention computation, to enable both bulky KV cache offloading and layer-wise KV cache offloading.
 The remainder of this section lists all the interfaces in Table 2, discusses the design for important APIs, and then traces how a query interacts with these interfaces end-to-end.
@@
 Putting it together, when a query comes in, the scheduler first calls get_num_new_matched_tokens which queries LMCache to see cache hit tokens in the backend.
+The function can return None if LMCache decides to let vLLM put the current request back to waiting queue and process other requests first, overlapping this request’s I/O with other requests’ computation.
 Then the update_state_after_alloc function decides whether each page in vLLM needs to be loaded from external storage backend based on matched tokens information from LMCache.
@@
 Then after the LLM inference has done for the current scheduling iteration, start_store_kv is called store the generated KV cache to lower-tier storage synchronously.
-Interface Description
-lookup(tokens) $\rightarrow$ {instance_id: hit_tokens} Returns the instance names of KV caches containing prefix-match tokens with the given token list.
-query_ip(instance_ids) $\rightarrow$ IP Returns the IP addresses for the corresponding instance_ids
-move(source, destination, tokens) Moves the KV cache of the given token list from a source instance to destination.
-clear(tokens, instance_id, storage_device) Clears the KV cache for corresponding tokens from the storage_device in instance.
-pin(tokens, instance, storage_device) Pins the KV cache for corresponding tokens at the storage_device in instance.
-compress(tokens, instance, storage_device, compression_method) Compresses the KV cache for the corresponding tokens with a compression method and store it at the storage_device in instance.
-Table 3: Functions in LMCache’s connector.
+Impact:  This API is out for over six months in vLLM.
+During these six months, we see open-source adoptions, including NVIDIA dynamo project, llm-d project from RedHat, AIBrix project from ByteDance, and vLLM production stack project.
+We also see multiple proprietary connectors from different companies that use the KV connector API.
+Internal APIs Description
+batched_admit/batched_evict(hashes, inst_id, device) Send the KV admission/eviction messages from an LMCache instance to the controller manager.
+batched_p2p_lookup(hashes) $\rightarrow$ list[inst_id, device, hit_chunks] Lookup peer KV cache existence from an LMCache instance based on the the given hashes.
+External APIs Description
+lookup(tokens) $\rightarrow$ list[inst_id, device, hit_tokens] Lookup the global KV cache existence of the given tokens.
+move((src_inst_id, src_device), (dst_inst_id, dst_device), tokens) Moves the KV cache of the given tokens from source location (src_inst_id, src_device) to destination location (dst_inst_id, dst_device).
+clear(tokens, inst_id, device) Clears the KV cache for corresponding tokens from the storage device device in instance inst_id.
+pin/unpin(tokens, instance, storage_device) Pins/unpins the KV cache for corresponding tokens at location (inst_id, device).
+compress/decompress(tokens, instance, device, method) Compresses/decompresses the KV cache for the corresponding tokens at location (inst_id, device) with a specified compression/decompression method.
+Table 3: APIs in LMCache Controller.
 7 Controller Interfaces
-There are many scenarios where higher-level applications need to see KV cache locations and operate on that.
-For example, query routing can benefit from knowing where the largest number of prefix cache hit tokens is at.
-To support such use cases, LMCache exposes APIs, as shown in Table 3, for querying and manipulating KV cache, such as locating cached entries or migrating them across instances.
-The KV cache controller in LMCache consists of two components: a centralized manager and per-instance workers.
-Incoming controller queries are handled by the centralized manager, which dispatches the appropriate operations to workers running within each LMCache instance.
-The remainder of this section illustrates how to use the controller interfaces through several example applications.
-KV cache–aware routing:  In this case, higher-level query routers would like to direct queries to instances with the largest number of prefix cache hits.
-The router first calls lookup(tokens), where the centralized manager dispatches the lookup queries to individual LMCache instances, to identify the instances containing matching prefix tokens, then invokes query_ip(instance_ids) to map instance ids to IP addresses.
-Finally, the query is forwarded to the IP with the highest number of hits.
-KV cache migration:  When an instance holding KV cache fails or load balancing is required, the KV cache may need to be migrated across instances.
-The move(source, destination, tokens) interface moves the KV cache corresponding to tokens from the source instance to the destination.
+LMCache operates as a distributed caching system built around a centralized KV cache controller responsible for global metadata management, cache manipulation, and request routing.
+To support these functionalities, LMCache provides two categories of APIs: (1) external APIs, which are directly accessible to users or system operators, and (2) internal APIs, which are used by individual LMCache instances.
+Mechanically, the KV cache controller consists of two layers: a centralized controller manager and per-instance workers.
+The controller manager runs as a standalone process and serves as a global coordination point, while per-instance workers collocate with each peer LMCache instance and handle local operations or issue global requests to the manager.
+External API calls are handled by the centralized manager, which, if necessary, dispatches the appropriate operations to each worker.
+Per-instance workers can also proactively interact with the centrailzed manager via internal APIs for metadata update or lookup.
+The KV cache controller underpins a series of advanced optimizations, including cross-node KV cache sharing, cache-aware request routing, and dynamic KV cache migration.
+The remainder of this section demonstrates how these optimizations could leverage the controller interfaces through concrete examples.
+KV cache–aware routing:  In this case, higher-level routers aim to direct requests to the instance with the highest expected cache hit rate.
+Each LMCache instance reports its cache admission and eviction decisions to the controller manager via the batched_admit and batched_evict interfaces.
+The controller manager aggregates these updates and maintains a global in-memory view of the KV cache state across all instances.
+When the router calls lookup(tokens), the controller consults its in-memory global KV cache states and returns a list of (instance_id, storage_device, hit_tokens), indicating where and how many of the requested tokens are currently cached.
+KV cache migration:  When an instance holding KV cache is about to be scaled down or load balancing is required, the KV cache may need to be migrated to another instance.
+The controller manager hadles such operations through move((src_inst_id, src_deivce), (dst_inst_id, dst_deivce), tokens) API call by dispatching the request to the source instance.
+The source instance will try to establish connection to the destination instance if one does not exist, and transfers the specified KV cache from the source storage device src_device to the destination location indicated by (dst_inst_id, dst_deivce).
+P2P KV cache sharing:  LMCache supports peer-to-peer KV cache sharing, allowing an instance to fetch KV cache from another peer when a local cache miss occurs.
+Upon a cache miss, the instance’s local worker can query the centralized controller manager via batched_p2p_lookup.
+The manager will return a list of (inst_id, device, hit_chunks), representing the number of hit chunks and the location that hold these chunks.
+The instance can then choose to, for example, load KV cache from a peer with maximum hit_chunks.
 KV cache clearance:  Applications may clear cache when switching models or reclaiming memory.
-The clear(tokens, instance, location) function removes the KV cache corresponding to tokens stored at a specific storage location (e.g., GPU memory or CPU memory) within the given instance.
-Pinning KV cache in GPU memory:  Some contexts, such as system prompts in chatbot applications, should remain in GPU memory for fast access.
-The pin(instance, location, tokens) function pins the specified KV cache in the chosen storage tier (GPU memory) of the given instance.
-There are other more advanced interfaces users can call to manipulate the KV cache, such as the compress(tokens, instance, location, compression_method) function which compresses KV cache stored in a specific location in a LMCache instance with compression_method so that the KV cache takes smaller amount of space in storage.
-Users can freely call these interfaces to explicitly manage the KV cache in their own applications based on their needs.
+Upon recieving a clear(tokens, inst_id, location) call, the controller manager dispatches the operation to the corresponding instance identified by inst_id.
+The instance’s worker then removes the KV cache associated with tokens stored from a specific stroage device device.
+Some APIs are not covered in the above applications, such as compress/decompress(tokens, inst_id, device, compression_method) which compresses/decompresses KV cache stored in location (inst_id, device) with a specified compression_method and pin/unpin(tokens, inst_id, device) which can pin/unpin the specifed KV cache at a certain location (inst_id, device).
+Users can freely call these APIs to explicitly manage the KV cache in their own applications based on their needs.
 8 Evaluation
-Figure 5: Single-node evaluation results.
-Figure 6: Real-trace evaluation results.
-Figure 7: Remote backend evaluation results.
-Figure 8: TTFT and ITL CDF comparison for PD.
-Figure 9: Comparing to SGLang CPU Offloading result.
-Figure 10: With request asynchronization, LMCache overlaps KV cache loading and inference computation (either prefill or decode).
-Figure 11: Latency breakdown of PD.
+[…]
 8.1 Setup
-Scenario Acronym
-Single-node /
-Multi-node
-Network
-Medium
-Real-world
-Examples
-CPU Offload Single-node N.A.
-Single-node
-CPU Offloading
-Central Storage Single-node Ethernet
-Centralized
-Storage Server
-PD Single-node NVLink
-PD
-Disaggregation
+Scenario Acronym Single-node / Multi-node Network Medium Real-world Examples
+CPU Offload Single-node N.A. Single-node CPU Offloading
+Central Storage Single-node Ethernet Centralized Storage Server
+PD Single-node NVLink PD Disaggregation
 Table 4: Evaluation scenarios setup.
 We evaluate LMCache under three different scenarios, as shown in Table 4. The three scenarios are representative setups that are commonly used by the users of LMCache.
-Models:  We compare LMCache against baseline solutions on popular open source models adopted by industry: meta-llama/Llama-3.1-8B-Instruct, meta-llama/Llama-3.1-70B-Instruct, Qwen/Qwen2.5-Coder-32B-Instruct, Qwen/Qwen3-Coder-480B-A35B-Instruct-FP8, Qwen/Qwen2.5-72B-Instruct.
-Datasets:  LMCache is evaluated on several datasets, including emulated multi-round question answering, long context question answering from LongBench Bai et al.
-(2024), and random dataset from vLLM official benchmarking script Kwon et al.
-(2023c).
+Models:  We compare LMCache against baseline solutions on popular open source models adopted by industry: meta-llama/Llama-3.1-8B-Instruct, Sao10K-L3-8B, meta-llama/Llama-3.1-70B-Instruct, Qwen/Qwen2.5-Coder-32B-Instruct, Qwen/Qwen3-Coder-480B-A35B-Instruct-FP8, Qwen/Qwen2.5-72B-Instruct.
+Datasets:  LMCache is evaluated on several datasets, including emulated multi-round question answering, long context question answering from LongBench bai2024longbenchbilingualmultitaskbenchmark, and random dataset from vLLM official benchmarking script vllm.
 […]
@@
 For component-wise analysis which breaks down the delay for CPU offloading or PD disaggregation, we report the delay for each component separately.
-Baselines:  We compare LMCache with several baselines, including 1) vLLM with prefix caching, which implements prefix caching, but only keeps limited amount of KV cache inside GPU memory; 2) commercial offering A, B, and C, which provide dedicated endpoint service that reserves GPUs for users to run a user-defined model.
+Baselines:  We compare LMCache v0.3.6 with several baselines, including:
+  • $\bullet$
+    Basic vLLM: vLLM v0.10.2 which enables prefix caching by default, but only keeps KV cache inside GPU memory, so only a small portion of it can be kept;
+  • $\bullet$
+    Basic vLLM CPU Offloading: vLLM v0.11.0 with its own implementation of CPU offloading;
+  • $\bullet$
+    Commercial offerings #1 and #2: provides dedicated endpoint service that reserves GPUs for users to run a user-defined model. We ran these baselines accessed on September 10th.
 8.2 Single-node CPU Offloading
@@
 Understanding LMCache’s gains:  LMCache outperforms baselines for several reasons.
-Compared with naïve vLLM prefill, which caches KV data only in GPU memory, LMCache leverages CPU offloading.
+Compared with basic vLLM, which caches KV data only in GPU memory, LMCache leverages CPU offloading.
 Since CPU memory can hold far more KV cache than GPU memory, LMCache achieves significantly higher cache hit ratios.
-With its efficient CPU-to-GPU loading scheme, cached data can be fetched quickly to accelerate inference.
+Compared to basic vLLM CPU offloading, which employs per-layer and per-16-token transmission and is unable to fully utilize the loading bandwidth, LMCache has more efficient implementation of the transmission module as it loads KV cache at chunk level with high-performance data loading CUDA kernels.
 Our comparison with closed-source commercial alternatives is conducted in a black-box manner since their internal implementations are not publicly available.
@@
 In contrast, Commercial Option #2 likely supports KV cache offloading to secondary storage, yet its performance is still worse than LMCache.
+8.3 Real-trace Driven Evaluation
+[…]
+[…]
+We evaluate LMCache on a real trace drawn from company F and company G’s distribution of input and output tokens.
+Since we do not have access to company $F$’s proprietary models, we run the trace using five different models.
+To make the experiment tractable, we stretch the original trace which lasts for several days so that the workload we run completes within one hour.
+LMCache is set to use 500 GB of CPU DRAM at maximum, and we compare it with latest basic vLLM with GPU prefix caching.
+[…]
 8.4 Centralized Storage Server
+[…]
 Next, we run LMCache for KV cache sharing through a centralized remote server, that is connected to the GPU instance with a bandwidth of 15 Gbps, following the setup of central storage in Table 4.
-For this experiment, we evaluate using the TriviaQA dataset from LongBench Bai et al.
-(2024), a widely adopted benchmark for long-context evaluation.
-We follow the official vLLM benchmarking scripts Kwon et al.
-(2023c), which generate inference queries according to a Poisson distribution at a specified QPS.
+For this experiment, we evaluate using the TriviaQA dataset from LongBench bai2024longbenchbilingualmultitaskbenchmark, a widely adopted benchmark for long-context evaluation.
+We follow the official vLLM benchmarking scripts vllm, which generate inference queries according to a Poisson distribution at a specified QPS.
 […]
 We note, however, that loading KV cache from the remote backend introduces greater latency than loading from CPU memory, since the remote backend has a much lower bandwidth.
-As a result, the loading delay may even surpass the prefill delay, particularly when the input context is short or model is small, as prefilling is too fast in such cases—a scenario that we will demonstrate later in 7.7.
+As a result, the loading delay may even surpass the prefill delay, particularly when the input context is short or model is small, as prefilling is too fast in such cases—a scenario that we will demonstrate later in 8.7.
+Thus, adaptive decisions between KV cache loading and prefilling need to be made when KV cache resides in a remote storage server.
 8.5 PD disaggregation
+[…]
+Figure 13: With request asynchronization, LMCache overlaps KV cache loading and inference computation (either prefill or decode).
+Figure 14: Compared to vLLM’s native PD disaggregation, LMCache achieves much smaller transmission latency, thus reducing end-to-end delay.
 […]
@@
 However, when the paged memory for the KV cache is scattered across the prefiller’s GPU memory, the transfer is performed in a page by page manner, which leads to bandwidth underutilization, as discussed in §5.
-7.5 Real-trace Driven Evaluation
-We also evaluate LMCache on a real production trace provided by company FF [^4] Anonymized due to NDA..
-The trace includes query arrival timestamps and actual multi-round chat inputs.
-Notably, later rounds within the same user session often reuse conversation history from earlier rounds, resulting in prefix reuse.
-On average, the inputs in the trace are about 4K tokens long.
-Since we do not have access to company $F$’s proprietary models, we replay the trace using the Sao10K/L3-8B-Lunaris-v1 model.
-To make the experiment tractable, we stretch the original trace which lasts for several days and apply random sampling so that the workload we run completes within one hour.
-As shown in Figure 6, LMCache consistently outperforms naive vLLM on the real trace on different QPS, reducing both TTFT and ITL, demonstrating its effectiveness in real-world settings.
-Note that for lower QPS from 2 to 5, LMCache still outperforms naive vLLM by 25%, while when QPS is 6, LMCache has improvement of 49% compared to naive vLLM.
 8.6 Component-wise Evaluation
@@
 We also perform several sensitivity evaluation to see how LMCache’s delay changes under different context lengths and different types of remote backends.
-Figure 12: Impact of different context lengths.
+Figure 15: At network bandwidth of 32Gbps, LMCache’s KV cache offloading only outperforms basic vLLM’s prefill when input length is more than 256K tokens.
+At network bandwidth of 64 or 128Gbps, LMCache’s KV cache offloading is better than prefill across all input lengths.
 Impact of context lengths:  Figure 15 shows the prefill delay on B200 machines and the loading delay under different network bandwidths.
@@
 8.8 SGLang Results
+Figure 16: On Qwen3-32B model, LMCache’s CPU offloading achieves comparable performance as SGLang’s native CPU offloading.
 Although our primary evaluation uses vLLM, we also evaluate LMCache integrated with SGLang.
@@
 9 Real-World Lessons and Experience
-8.1 Scaling KV Cache Storage and Offloading
-One clear trend in the adoption of LMCache in production systems is that more and more KV cache needs to be stored, definitely beyond the number that can be stored in GPU memory.
-Early deployments kept all KV caches in GPU memory, but long contexts and multi-user workloads may quickly exhaust GPU memory.
-Thus, many companies now need to offload KV data to CPU RAM, CPU memory pooling, or even disks when needed.
-This prevents evicting useful cache entries — without offloading, a follow-up query would trigger full prefill of the same input context, often long contexts, which greatly prolongs TTFT.
-Interestingly, some teams, including company R, have experimented with remote storage devices.
-A remote data fetch is orders of magnitude slower than GPU memory access, yet real-world tests have shown it can still yield both speedups and cost savings.
-Retrieving previously computed KV cache over the network can be faster than recomputing them on loaded GPUs, with LMCache’s KV cache loading optimizations (§4).
-While the overhead of loading KV cache from remote is large, it can be pipelined with inference computations, either prefilling requests that do not have cache-hit tokens or decoding computations.
-These experiences show that even slow and cheap storage can be leveraged to efficiently store and load KV caches.
-Another technique used in production is the combination of CPU offloading and PD disaggregation.
-At company T, on the prefiller side, the KV cache is not only sent to the decoder side, but is also offloaded to the CPU memory of prefiller instance to achieve both the benefit of CPU offloading and PD disaggregation.
-8.2 Emerging Use Cases and Surprising Outcomes
-While chatbots were one of the first obvious beneficial usage scenarios of KV cache reuse, new use cases are rapidly emerging.
-Recently, large-scale recommendation systems from company L and company M have also adopted LLMs as the recommender model and becomes an important use case of KV cache reuse.
-In these systems, LLMs are used as the embedding models to prefill the contexts, without generating tokens, and are frequently reused by different queries of the same user.
-Furthermore, the contexts are often very long.
-Caching those KV cache can directly eliminate the expensive prefill phase, thus is a perfect use case of LMCache.
-Although multi-round chat has long been a canonical use case for KV cache reuse, real-world deployments have revealed several important insights.
-Many systems initially adopted sliding context windows to fit more tokens into limited GPU memory during inference.
-However, this approach often degrades generation quality by discarding useful history and reduces prefix cache hit ratio, since truncated contexts are no longer prefixes.
-Deployment experiences at Company T and F demonstrate that keeping the full conversation history, while offloading KV caches to larger memory tiers, is a more effective practice.
-By avoiding truncation, the system achieves significantly higher cache hit rates on shared prefixes, leading to less GPU memory usage per request.
-Moreover, in production environments, companies have observed that prefix cache hit rate is far higher than they expected.
-The experience from Company B’s adoption shows that industry users are accepting lossy optimizations, such as KV cache compression, for better system performance.
-Especially, in open-ended chatbot applications, where there is no single “correct” answer, using compressed KV cache (e.g., quantized KV cache) may not harm the user experience but can dramatically reduce memory usage and I/O time.
-Even in finance companies, teams found that slight differences from using a compressed KV cache are tolerable if the overall system throughput improves.
-8.3 Integration and Deployment Challenges
-Deploying a caching layer in production LLM inference systems has its own practical challenges and lessons.
-One key lesson we found in the adoption of LMCache is that many companies preferred containerized environments, such as Docker images, when deploying the caching layer into production, as they often do not go into the source code of LMCache.
-Another important aspect in the caching layer for LLM inference is fault tolerance and transparency.
-Because the caching layer sits beneath the inference engines and is supposed to be invisible to end users, robustness is important.
-Many companies stressed that if the cache system fails or underperforms, it should not bring down the model service.
-For instance, LMCache incorporates fault-tolerant KV cache retrieval: if errors occur during retrieval, LMCache reports which portions of the KV cache were successfully read, ensuring the inference engine does not break due to the unsuccessful read of KV cache and the correctness of generation is preserved.
-We also observed a growing demand for a clean separation between the model serving engine and the caching engine.
-Many storage companies, such as company I, R, C, and W, have explicitly asked for decoupling the KV cache management and the core inference code, since they want to touch as minimum core inference logic as possible.
-Although LMCache does not solve this for now, the trend is clear: an LLM inference engine is pairing with a variety of external KV cache management services.
-At last, we found that ensuring compatibility between LMCache and emerging inference frameworks was easier than expected.
-LMCache was originally designed and developed on top of vLLM.
-However, integrating LMCache with other frameworks, such as SGLang, has also been straightforward, as long as they adopt the paged attention mechanism.
-Looking ahead, we expect that LMCache will be equally easy to integrate with other inference frameworks based on paged attention.
-8.4 Adoption, Lessons Learned, and Future Outlook
-Our journey began as an academic prototype in the summer of 2024.
-At that time, KV caching was a niche concept mostly discussed in a few research papers.
-But by spring 2025, the landscape had changed — KV cache became a key technique in industry for improving the efficiency in LLM inference.
-The adoption of LMCache truly exploded in early 2025, with numerous companies starting to collaborate and use.
-Interestingly, although LMCache has initially been an academic prototype, we found that LMCache has much more industry users than academia ones.
-One potential reason is that there is lack of an easy set of programmable interfaces.
-As a concrete example, a rich set of research literatures explore dropping tokens from KV cache for compressing it into smaller tensors.
-However, it is non-trivial to implement token dropping in LMCache, since LMCache assumes the number of tokens that is input to and output of LMCache should be the same.
-Furthermore, any intermediate states in the attention module are hidden from the LMCache’s side, so research works that need to access or modify the attention states are hard to be implemented with LMCache.
-Finally, we learned that programming languages for high performance are not always the best fit for building LLM inference infrastructure.
-Early on, some organizations (e.g.,, company R and I) assumed that KV cache management must be implemented in Rust or C++ to achieve high efficiency.
-In contrast, we developed LMCache in Python, which turned out to be a more effective choice.
-Python not only simplifies integration with existing inference frameworks but also lowers the barrier for community contributions, enabling rapid feature development and iterations.
-That said, CUDA-based implementations are still essential for certain high-performance data loading modules within LMCache.
+Loading from remote storage is faster than prefill:  Traditionally, it was believed that loading KV cache from remote storage primarily improves cache hit rates and reduces storage costs by leveraging cheaper storage devices, but at the cost of increased inference latency, since loading data from remote devices was thought to be slower than performing a full prefill.
+This assumption was largely due to the historically low throughput of remote object stores such as Amazon S3, which offered loading speeds as low as 100 MBps.
+Recently, however, there has been a major boost in remote storage performance, for example, with Amazon S3 Express, the throughput has increased from 100 MBps to nearly 1 GBps.
+Our users, such as Company C, have adopted LMCache to load KV caches from their own remote object store, achieving 22–32% lower TTFT compared to full prefill.
+This insight suggests that remote backends can simultaneously improve cache hit ratios and reduce TTFT.
+Context truncation lowers prefix cache hit rates:  Many industry users employ a sliding-window mechanism to handle long-context inputs constrained by limited model context windows or GPU memory.
+For instance, when input tokens exceed the context window limit, some companies truncate the input to keep only the most recent tokens.
+However, this approach significantly reduces prefix cache hit ratios, since truncated inputs no longer match the prefixes of previously cached contexts.
+In practice, using real traces from Company F, we observe that prefix cache hit ratios drop from roughly 85% to 45% when truncating input contexts to keep only the latest tokens.
+Other studies have also discussed this phenomenon, emphasizing that dynamically adding or removing context tokens should be avoided, as it invalidates prefix KV cache reuse manus_blog.
+Containerized code is preferred:  With the growing scale of LLM inference, most production environments now rely on Kubernetes to manage GPU clusters.
+Consequently, deploying inference engines (e.g., vLLM or SGLang) and LMCache through containerized environments, typically via Docker images, has become the standard practice among industry users.
+Interestingly, many users rely solely on the official Docker images without going deep into LMCache ’s source code.
+Unexpected high cache hit rate in production systems:  Our customers did not expect such a high prefix cache hit ratio, such as 50% hit rate for company G in their production environments, until they deploy LMCache inside their systems.
+Previously, people thought KV caches could only be reused for fixed system prompts.
+However, modern applications increasingly exhibit "dynamically reusable contexts", such as conversation histories in coding assistants, chat applications, and retrieval-augmented generation (RAG) pipelines.
+These emerging patterns have significantly increased overall cache hit ratios in real-world deployments.
+Industry vs.
+academia users:  LMCache was designed as a unified prototype framework in May 2024, where we can put our research works into it to gain more impact.
+However, we found that at that time, industry wants a highly efficient KV cache offloading solution as the size of KV cache and concurrent users keep growing.
+Then, our attention has been moved towards improving the performance, stability, and compatibility.
+Since most companies are less concerned with customizing attention algorithms, we deprioritized designing flexible APIs for integrating specialized attention mechanisms, such as selective token dropping.
+This makes LMCache less popular in academia, since the academia users often focuses on modifying attention mechanisms for their research prototypes.
+As next step, LMCache is going to design more flexible APIs such that it is easy to use by both industry and academia.
+Flexibility vs.
+performance for programming language:  Python has always been the de facto language in ML.
+However, the current industry focus is gradually shifting towards higher efficiency instead of broad compatibility.
+Specifically, many companies are rewriting ML libraries in high-performance languages such as Rust or C++, or carefully optimizing Python-based systems to hide its runtime overhead while preserving flexibility.
+Although some alternatives have been writing their ML libraries in Rust, we continue to use Python with carefully designed optimizations.
+This approach allows us to evolve faster, with more community contributions, and still maintain performance on par with alternatives.
+LMCache now a community effort:  One key reason LMCache rapidly evolved from a research prototype to a widely adopted industry framework is the active involvement of community contributors.
+About a year ago, LMCache supported only local CPU, local disk, and Redis backends integrated with vLLM on NVIDIA GPUs.
+Today, it supports eight more storage backends (NFS, WEKA, GPU-Direct Storage, Mooncake Store, NIXL, S3, InfiniStore, and Valkey) across four processor types (NVIDIA, AMD, Ascend, and TPU), and two inference engines (vLLM and SGLang).
+All of these contributions were made by industry partners who actively upstreamed their code to stay aligned with ongoing development and avoid divergence from the latest LMCache updates.
 10 Conclusion and Outlook
@@
 11 Acknowledgement
-We would like to thank the LMCache community for their invaluable support and contributions, including Baolong Mao and Chunxiao Zheng for managing remote connectors, Martin Hickey for GitHub Infrastructure, Huaizheng Zhang, Siddhant Ray, Zhuohan Gu and Hanchen Li for writing and maintaining documentation, Rui Zhang for the deployment of LMCache, Qizheng Zhang and Hussain Mohammad for insightful feedback.
+We would like to thank the LMCache community for their invaluable support and contributions, including Baolong Mao and Chunxiao Zheng for managing remote connectors, Martin Hickey for GitHub Infrastructure, Huaizheng Zhang, Siddhant Ray, Zhuohan Gu and Hanchen Li for writing and maintaining documentation, Qizheng Zhang and Hussain Mohammad for insightful feedback.
+We also thank GMI cloud for providing us GPU clusters to run the experiments.